### v1

# LocMoE+: Enhanced Router with Token Feature Awareness for Efficient LLM Pre-Training

Abstract
Mixture-of-Experts (MoE) architectures have recently gained increasing popularity within the domain of large language models (LLMs) due to their ability to significantly reduce training and inference overhead. However, MoE architectures face challenges, such as significant disparities in the number of tokens assigned to each expert and a tendency toward homogenization among experts, which adversely affects the model’s semantic generation capabilities. In this paper, we introduce LocMoE+, a refined version of the low-overhead LocMoE, incorporating the following enhancements: (1) Quantification and definition of the affinity between experts and tokens. (2) Implementation of a global-level adaptive routing strategy to rearrange tokens based on their affinity scores. (3) Reestimation of the lower bound for expert capacity, which has been shown to progressively decrease as the token feature distribution evolves. Experimental results demonstrate that, without compromising model convergence or efficacy, the number of tokens each expert processes can be reduced by over 60%. Combined with communication optimizations, this leads to an average improvement in training efficiency ranging from 5.4% to 46.6%. After fine-tuning, LocMoE+ exhibits a performance improvement of 9.7% to 14.1% across the GDAD, C-Eval, and TeleQnA datasets.
1 Introduction
Large language models (LLMs) have shown exceptional proficiency in understanding deep structures and complex semantic relationships within language [1]. As these models scale up, their capabilities in language generation and logical comprehension are enhanced, but this comes at the cost of significant computational, communication, and storage demands [2]. To scale models efficiently without disproportionately increasing computational costs, researchers have incorporated the Mixture-of-Experts (MoE) architecture into LLMs [3]. The MoE framework integrates multiple experts within the model, each tasked with processing specific types of inputs [4]. For a given input, only a subset of experts is activated, allowing for more efficient use of computational resources [5]. Recently, several LLMs employing MoE structures, such as DeepSeek-MoE [6] and Mixtral [7], have demonstrated outstanding performance on various leaderboards.
Despite the efficiency benefits of MoE in scaling model sizes, it introduces several new challenges and drawbacks [8]. For instance, an imbalanced "winner-takes-all" phenomenon among experts can detrimentally affect overall model performance [9], and the All-to-All communication pattern can become a bottleneck [10]. The model’s convergence and the experts’ generalization capabilities are heavily dependent on the design of the routing strategy. Recent research has addressed these challenges from multiple angles, such as optimizing routing strategies and enhancing communication efficiency. In terms of routing strategies, DeepSeek-MoE [6] introduced the concepts of vertical-class experts, created through fine-grained partitioning, and shared experts, which are always activated. Google Brain proposed the Expert Choice (EC) routing algorithm [11], assigning experts with predetermined buffer capacities to the Top-k tokens to ensure load balance. For communication efficiency, FastMoE [12] leverages Megatron-LM [13], a distributed training framework by NVIDIA, utilizing CUDA’s multi-stream mechanism to maximize the overlap between communication and computation. Tutel [14] introduced flexible All-to-All communication to alleviate throughput bottlenecks caused by smaller matrix dimensions and proposed 2DH All-to-All to align non-contiguous memory spaces and merge scattered communications. The uneven expert load in classical MoE [15] often necessitates token dropping or padding. Megablocks [16] employs block-sparse operations to overcome the limitations of dynamic routing, improving the balance between model quality and efficiency.
LocMoE [17] aims to enhance the interpretability of routing mechanisms while reducing computational and communication overhead. LocMoE leverages orthogonal routing weights to prevent token homogenization across different expert networks and introduces the Grouped Average Pooling (GrAP) layer [18] for token feature extraction. Under these conditions, LocMoE uses theoretical proofs to calculate the lower bound of expert capacity. Experimental results demonstrate that employing this lower bound does not degrade, but slightly enhances, performance on downstream tasks. Additionally, LocMoE introduces a loss regularization term that incorporates data locality and load balancing [19], leading to a 12% to 22% improvement in training efficiency.
Building on previous work, this paper proposes LocMoE+, which aims to further enhance training efficiency and model performance. LocMoE+ adopts GrAP and the locality loss regularization terms from LocMoE, and continues to analyze the impact of token features on expert selection. The primary contributions of this paper are as follows:
  1. 1.
    The affinity score. Given that expert selection is related to the angle between tokens and gating weights, the cosine similarity between tokens and gating weights is used to define the affinity score between tokens and experts. This score is utilized for token rearrangement and as a threshold for the lower bound of expert capacity.
  2. 2.
    The hybrid token rearrangement scheme. By integrating the concepts of Token Choice Router (TCR) and Expert Choice Router (ECR), the training success rate is enhanced while considering expert capacity constraints. Its effectiveness has been theoretically validated.
  3. 3.
    The adaptive expert capacity. Setting a holistic affinity threshold allows the lower bound of expert capacity to be significantly reduced. As training iterations increase, the information density of token features grows, causing the expert capacity to initially decrease and then stabilize, thereby enhancing training efficiency.
LocMoE+ uses the state-of-the-art MoE model Mixtral 8$\times$7B as the backbone, and utlizes AscendSpeed and ModelLink libraries for pre-training on Ascend NPU clusters. Experiments conducted on clusters with 32, 64, and 256 NPUs show that LocMoE+ improves training efficiency by 5.4% to 46.6% compared to the baseline, and by 2.9% to 13.3% compared to LocMoE. Model performance is enhanced by 9.7% to 14.1% compared to the baseline, and by 1.7% to 4.1% compared to LocMoE.
The rest of this paper is structured as follows: Section 2 introduces related work on LLMs, MoE, and the Ascend architecture. Section 3 reviews LocMoE and presents the methods proposed in this paper, along with theoretical evidence. Section 4 analyzes the experimental results of LocMoE+ regarding training efficiency and model performance. The final section summarizes the content of this paper and offers an outlook on future improvements.
2 Related Work
LLMs.
BERT [20] and GPT [21] in 2018 marked the emergence of the LLM era. GPT-3 [22], with 175 billion parameters, demonstrated impressive capabilities in language understanding and generation in 2020. By 2023-2024, LLMs expanded to hundreds of billions to trillions of parameters: GPT-4 [23] achieved breakthroughs in zero-shot and few-shot learning; PaLM-2 [24] enhanced knowledge acquisition and reasoning; Chinchilla [25] optimized computational resources and model scale; and Claude-v1 [26] displayed a safer, human-aligned LLM. These LLMs have made significant advancements across domains, paving the way towards artificial general intelligence.
MoE.
Despite the outstanding performance of LLMs in language understanding and generation, their high computational cost hinders further model scale expansion. To alleviate this issue, researchers have introduced the MoE architecture, which incorporates multiple expert modules and dynamically selects a subset of experts to process inputs, significantly reducing computational overhead while maintaining performance. GShard [3] and Switch Transformer [4] employ Top-2 and Top-1 routing strategies, respectively, to mitigate the expert load imbalance problem. Hash Layer [27] adopt hashing techniques to simplify the gating mechanism and improve model performance. METRO [28] introduces a metric learning-based routing mechanism that guides expert selection by learning task-relevant metric functions, enhancing the model’s generalization ability. DSelect-k [29] proposes a dynamic selection algorithm based on Markov chains, adaptively choosing the number of experts to maintain performance while reducing training costs. Furthermore, DeepSpeed-MoE [30] optimizes expert parallelism and pipeline parallelism in MoE, further boosting training efficiency. These works lay the foundation for the subsequent application of MoE in LLMs.
Ascend Architecture.
The AI Core in the Ascend 910B NPU adopts the DaVinci architecture [31], which is responsible for executing computation-intensive operators related to scalars, vectors, and 3D cubes, and is widely used in LLM training. Huawei’s proprietary cache coherence protocol interconnect, High Confidence Computing Systems (HCCS) [32], enables high-performance inter-device data communication in multi-card scenarios. The Huawei Collective Communication Library (HCCL) [33], designed for the Ascend processor, offers high-performance primitives for single-node multi-card and multi-node multi-card collective communication. It implements collective communication functions on PCIe, HCCS, and RoCE high-speed links, facilitating distributed training. PyTorch [34], a widely adopted deep learning framework, is compatible with the Ascend NPU through the Ascend extension for PyTorch [35]. AscendSpeed, an acceleration library tailored for Ascend devices, focuses on parallel strategy partitioning for LLMs and aggregates results through collective communication. ModelLink, a training framework designed for LLMs, fully considers the characteristics of the Ascend NPU and implements features such as memory optimization, multi-dimensional parallelism, and adaptive data loading, further enhancing training performance.
3 Method
Notation.
[…]
3.1 Background
Main Principle.
The MoE architecture, based on the Transformer framework, efficiently scales up model size with low computational overhead, benefiting from two primary structures: a sparse gating network for routing tokens and expert networks for processing specific token categories. The sparse gating network determines the expert to which a token will be sent.
We consider the supervised classification for brevity where the training samples are $\{({\bm{x}}^{(i)},y_{i})\}_{i=1}^{N}\sim{\mathcal{D}}$. Each training sample ${\bm{x}}^{\top}=({\bm{x}}_{1}^{\top},\dots,{\bm{x}}_{s}^{\top})\in{\mathbb{R}}^{sd}$ has $s$ tokens with token feature ${\bm{x}}_{i}\in{\mathbb{R}}^{d},\forall i\in[s]$, and label $y\in{\mathbb{N}}^{+}$. The objective is to learn the map of ${\bm{x}}$ to the corresponding $y$. The general MoE structure are formulated as
$\text{MoE}({\bm{x}})=\sum_{t=1}^{s}\sum_{i=1}^{n}G_{i}({\bm{x}}_{t})\cdot E_{i}({\bm{x}}_{t}),$ (1)
where $n$ is the number of experts, $G({\bm{x}}_{t})\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{n}$ is the gating (or router) weights of experts which maps the tokens of ${\bm{x}}_{t}$ into the coresponding experts with weights, e.g., $G_{i}({\bm{x}})=\text{Softmax}({\bm{W}}{\bm{x}}+\bm{\epsilon})$ where the softmax is applied to each row, and $E_{i}({\bm{x}}_{t})\colon{\mathbb{R}}^{d}\to{\mathbb{R}}$ is the $i$-th expert network, e.g., simple MLP models, see [36] for current different router methods. Generally, $n\ll s$, which saves much computation compared to the dense structure.
Routing Strategy.
There are two main router methods: hard router [36] which uses hard assignments between tokens and experts with sparse dispatch matrix, and soft router [37] which uses weighted combinations of tokens. In this paper, we still focus on the first one, since the computation cost is much smaller than the soft ones.
The hard router also have two main choices: 1) Token Choice Router (TCR) [38, 11], which lets each token choose its top-scored experts; Taking Top1 routing as an example, it typically selects a linear layer to extract feature representations for different experts and then obtains the expert index with the highest routing probability through the gating function. 2) Expert Choice Router (ECR), which lets each expert choose its top-scored tokens. Moreover, to balance the budget of differnet experts, we will pre-define the expert capacity $C$ of all experts in practice, namely, the maximum tokens sent to each expert.
Chocie of score. In this paper, we define the score of $t$-th token and $i$-th expert as $\delta_{ti}=\cos\left({\bm{x}}_{t},{\bm{w}}_{i}\right):={\bm{x}}_{t}^{\top}{\bm{w}}_{i}/(\|{\bm{x}}_{t}\|\cdot\|{\bm{w}}_{i}\|)$, i.e., the cosine simlarity between vectors ${\bm{x}}_{t}$ and ${\bm{w}}_{i}$. The score is to decide the gating weight in router. For example, we can route $t$-th token to the experts of the $-\ell$ indices in $(\delta_{t1},\dots,\delta_{tn})$.
Chocie of capacity. In previous works, expert capacity $C$ is calculated using a capacity factor, ensuring that the size of each expert’s token buffer remains consistent. The choice of capacity factor is currently based mainly on empirical knowledge; increasing the capacity factor allows for handling more tokens but reduces training efficiency.
TCR expression. First, given $\ell$, we decide the candidate expert set for each token by the sort of scores:
$\left(\tilde{I}_{t1},\dots,\tilde{I}_{t\ell}\right)=\text{Top-}\ell\left(\left\{\delta_{t1},\dots,\delta_{tn}\right\}\right),\tilde{I}_{tk}\in[n],\forall t\in[s],k\in[\ell].$ (2)
Second, we choose the first-$C$ tokens (or all if less than $C$) to each expert due to the capacity limit:
$\left(I_{1i},\dots,I_{Ci}\right)=\text{Bottom-}C\left(\left\{t\in[s]:\exists j\in[\ell],\tilde{I}_{tj}=i\right\}\right),I_{ki}\in[s]\cup\text{None},\forall i\in[n],k\in[C].$ (3)
In practice, we adopt $\ell=1$ to further save the computation cost since $n\ll s$.
ECR expression. We directly choose the tokens in the top-$C$ scores of each expert:
$\left(I_{1i},\dots,I_{Ci}\right)=\text{Top-}C\left(\left\{\delta_{1i},\dots,\delta_{si}\right\}\right),I_{ki}\in[s],\forall i\in[n],k\in[C].$ (4)
Intuitively, TCR can result in underused experts with fewer tokens than their allowed capacity, while ECR can result in underused tokens where no expert receives such token.
Auxiliary Loss.
Feed-forward network (FFN) layers are commonly employed in expert networks, allowing each expert to learn independently as a separate neural network, thus preventing interference between samples. This mechanism leads to a severe load imbalance, as experts frequently selected in the early stages are more likely to be chosen in later stages. To mitigate this skewness in token allocation, the auxiliary loss [8] has been proposed, defined as $L_{\mathrm{aux}}=\alpha{{n}{\sum_{i=1}^{n}f_{i}{P_{i}}}}$, where $f_{i}=\frac{1}{T}\sum_{{\bm{x}}\in{\beta}}\mathds{1}{\mathop{\arg\max}p({\bm{x}})=i}$, and $P_{i}=\frac{1}{T}\sum{{\bm{x}}\in{\beta}}p_{i}({\bm{x}})$. Here, $\beta$ represents the batch containing $T$ tokens, and $p_{i}$ denotes the probability of a token being routed to expert $i$.
3.2 LocMoE
GrAP.
[…]
The Lower Bound of the Expert Capacity.
Under the premises of orthogonal gating weights and a data distribution approaching uniformity, LocMoE demonstrates that the expert capacity is closely related to the angle between the gating weights and tokens. For large scale of the activation, the lower bound of expert capacity is proven to exist and is represented as $C_{\mathrm{min}}=\frac{1}{n}\exp\{d\delta_{\mathrm{max}}^{2}/(2-\delta_{\mathrm{max}}^{2})\}$.
Locality Loss.
Building upon the auxiliary loss, LocMoE introduces a loss bias term based on data locality, represented as $L_{\mathrm{loc}}=\mu\mathrm{KL}(D_{\mathrm{c}}||D_{\mathrm{l}})=-\mu\int{D_{\mathrm{c}}(x)\ln[\frac{D_{\mathrm{l}}(x)}{D_{\mathrm{c}}(x)}]\mathrm{d}x}$, i.e., the Kullback-Leibler (KL) divergence of the current distribution $D_{\mathrm{c}}(x)$ and the fully localized distribution $D_{\mathrm{l}}(x)$. This loss term serves as a soft constraint, encouraging tokens to be sent to experts residing on the same node, thereby mitigating the substantial overhead incurred by partial inter-node communication.
3.3 Token Awareness Router for LocMoE
3.3.1 Scheme Sketch
Figure 1 illustrates the model constructed in this paper, with a particular focus on the proposed token awareness router. The model framework is GPT-like, comprising only decoders, and incorporates LocMoE+ in place of the gated routing of sparse MoE blocks. To enhance interpretability and reduce computational overhead, the original Top-2 routing is simplified to the sparsest Top-1 routing. Consequently, the weights of GrAP, consisting of orthogonal vectors, emphasize different aspects of token features, which are evenly divided and non-overlapping. This routing mechanism can be viewed as a multi-classification of these feature fragments.
TCR, introduced in the background, is exemplified in the figure. If the feature fragment corresponding to the $k$-th dimension of the gating weight for a particular token is more prominent, then that token will be routed to the $k$-th expert. If among all tokens routed to the $k$-th expert, there is a certain probability of the presence of class-discriminative tokens, then the capacity $C$ must be set to a larger value to ensure the inclusion of sufficient class-discriminative tokens. The router proposed in this paper is a hybrid of TCR and ECR modes. After determining the expert to which a token will be routed, scores are calculated for the tokens assigned to each expert, and a $-\ell$ selection is performed, where $\ell^{*}$ is determined by a threshold of the sum of scores. Subsequent theoretical analysis will demonstrate the effectiveness of this hybrid routing scheme.
Figure 1: The architecture of the model along with the hybrid TCR+ECR router.
3.3.2 Theoretical Explanation
To explain the motivation of our method, we show some theoretical insights in this section. Our theoretical analyis is bulit on [39], where they assume the training data has unique class-discriminative patterns with exectly the same number as experts, and other class-irrelevant patterns. Chowdhury et al. [39] demonstrated that the classical MoE structure go through two phases:
Phase 1: Router training ([39, Lemma 4.1, Assumption 4.4]), which makes class-discriminative patterns all to the corresponding expert. This process ensures that each expert only receives the class-discriminative tokens related to the specific class.
Phase 2: Expert training ([39, Theorem 4.2, Theorme 4.5]), which makes each expert learn to predict the label based on its class-discriminative inputs from Phase 1. This process is designed to establish each expert’s ability to handle and solve problems.
Hence, the traning of an input in the current step is valid if the class-discriminative patterns is correctly dispatched. We compute training success rate of input (i.e., the probability of the tokens in an input is correctly dispatched to experts) to measure the difference between TCR and ECR,
We follow the setup in [39] that the input ${\bm{x}}\in{\mathbb{R}}^{sd}$ with $s$ tokens is comprised of one class-discriminative pattern ${\bm{o}}_{1},\dots,{\bm{o}}_{n}\in{\mathbb{R}}^{d}$, with each decides the label $y\in[n]$, and $s-1$ class-irrelevant patterns ${\bm{r}}\sim{\mathcal{N}}$ (the distribution will be clarified below). To show the quantitative comparison of TCR and ECR in training success rate, we need following asssumptions of token patterns.
Assumption 3.1 (class-discriminative).
We assume the location and feature of class-discriminative pattern is uniformly distribute in $[s]$ and $[n]$, i.e.,
$i\sim\mathrm{Unif}([s]),{\bm{x}}_{i}\sim\mathrm{Unif}\left(\{{\bm{o}}_{1},\dots,{\bm{o}}_{n}\}\right).$
We also assume that $\forall i\in[n],{\bm{o}}_{i}$ should be sent to the $i$-th expert, and define the true positive probability in token choice setting is no worse than the uniform dispatch as below
${\mathbb{P}}({\bm{o}}_{i}\text{ sent to expert }i)=p_{i}\geq 1/n,\forall i\in[n].$
Assumption 3.2 (class-irrelevant).
The distribution of class-irrelevant patterns is isotropy, i.e.,
${\mathbb{P}}({\bm{r}}\sim{\mathcal{N}},{\bm{r}}\text{ sent to expert }i)=1/n,\forall i\in[n].$ (5)
And we define the false positive probability in expert choice setting as
${\mathbb{P}}({\bm{r}}\sim{\mathcal{N}},\text{ the score of }{\bm{r}}\text{ is larger than }{\bm{o}}_{i}\text{ for expert }i)=q_{i},\forall i\in[n],$ (6)
which measures the possibility that expert $i$ chooses the wrong token ${\bm{r}}$ instead of the correct token ${\bm{o}}_{i}$.
Assumption 3.1 assumes the valid token is uniformly distributed in training samples due to the massive amounts of data nowadays. Assumption 3.2 assumes the invalid tokens can be uniformly dispatched to experts since the invalid tokens do not provide supervised signal to router and experts in training. We consider such uniform settings are common assumptions in theoretical analyis.
Now we compute the training success rate of TCR and ECR.
Theorem 3.3.
Under Assumptions 3.1 and 3.2, the training success rate of TCR in each sample ${\bm{x}}$ is
[…]
and the training success rate of ECR in each sample ${\bm{x}}$ is
[…]
Figure 2: The time consumption during training iterations with different schemes and cluster sizes.
Remark 3.4.
In practice, we have constant number of experts [7], i.e., $n=\Theta(1)$, and $C<s$ to save computation cost. We explain the benefit of swicthing TCR to ECR during training based on Theorem 3.3 and feature distrution during training.
At the beginning of training, the model seldom learn the task. Then the feature of class-irrelevant tokens is nearly isotropy, e.g., uniformly distrbute around the sphere (see Appendix B), leading to $q_{i}=\Theta(1)$. The succed rate of TCR with the form $C/s$ is better than ECR with the form $e^{-s}$. Thus we should choose TCR with a large capacity $C=\Theta(s)$ to improve the success rate of training samples.
After training for some iterations, the experts can roughly distinguish the class-irrelevant and discriminative patterns, leading to $q_{i}\ll 1$ or $sq_{i}\leq C^{*}$ for some $C^{*}>0$ (see Appendix B). Then ECR with success rate nearly $1$ is better than TCR with the form $C/s$ as long as $C\geq 2C^{*}$. Thus we should choose ECR with a small capacity $C=\Theta(1)$ to improve the success rate of training samples.
Indeed, we find that Chowdhury et al. [39, the definition of $\ell^{*}$] consider the ECR setting and verify the benefit in sample complexity. They assume the maximum number of class-irrelevant patches that are close to class-discriminative patches are bounded, which has similar effect as $C^{*}$ in our scene.
3.4 Communication Optimization
The training framework employs the Communication Over Computation (CoC) optimization technique to address performance bottlenecks in LLM training. During forward propagation in LLMs, the ColumnParallelLinear and RowParallelLinear components involve sequentially dependent computation (matrix multiplication) and communication (collective operations like AllReduce, AllGather, and ReduceScatter). These dependencies lead to inefficient serial execution. CoC decomposes these tasks into finer-grained subtasks and merges computation and communication into single kernels, such as MATMUL_ALL_REDUCE and MATMUL_REDUCE_SCATTER, utilizing MTE’s remote memory access capabilities. This approach allows for pipeline-style parallel execution and overlapping of computation and communication, significantly enhancing overall efficiency.
4 Experiments
4.1 Experimental Setup
The details of experimental setup including datasets, environment, configuration, and metrics, can be seen in Appendix C.
4.2 Efficiency Promotion and Memory Footprint Reduction
Figure 3: (a) The average composition of computation, communication, overlap, and idle with different schemes and cluster sizes. (b) The perplexity during training iterations with different schemes.
As detailed in Section 3, we consistently use Top-1 routing to ensure the routing implementation aligns with our theoretical framework. The Baseline model utilizes a limited expert capacity mode instead of the groupedGEMM scheme, which avoids token dropping, with the capacity factor set to 1.1. LocMoE considers data distribution uniformity and estimates expert capacity using a lower bound formula derived from its theoretical conclusions in the first batch, maintaining it as a constant during subsequent training. LocMoE+ fixes the range of score sums, processes hidden states, and calculates current expert capacity. The subsequent analysis addresses the training time, convergence, and memory usage efficiency of these schemes on multiple sizes of Ascend clusters.
Figure 2 illustrates the time consumption of these methods during the first 1000 iterations of training. Due to initialization and some unstable factors, time consumption is recorded starting from the 5th iteration. The Baseline model’s time consumption is relatively stable. As iterations increase, LocMoE’s time consumption slightly decreases, particularly in 32N and 64N, consistent with the conclusion that locality loss is effective only when the number of experts is greater than or equal to the number of nodes. LocMoE+ incurs slightly higher time consumption than LocMoE due to the computational overhead of token rearrangement. However, as token features converge, the required tokens gradually decrease and stabilize, leading to a decline in time consumption, which remains stable in subsequent training processes. Overall, LocMoE+ reduces training time by 2.9% to 13.3% compared to LocMoE, and by 5.4% to 46.6% compared to the Baseline.
We select 10 iterations at equal intervals from the training iterations to collect data on the time consumption of computation, communication, overlap, and idle periods, as shown in Figure 3(a). It is important to note that the data collection operation also introduces some overhead. After integrating LocMoE and LocMoE+, the time consumption of each component decreases, with a significantly greater reduction in computation overhead compared to communication overhead. Additionally, as the cluster size increases, the proportion of computation/communication overlap decreases, and the magnitude of the reduction in computation overhead diminishes. Figure 3(b) illustrates perplexity as a measure of convergence. The convergence curves of these approaches indicate normal loss convergence, with LocMoE+ not adversely impacting convergence.
The proportional time consumption at the operator level is depicted in Figure 4. Among the components, AI CORE efficiently executes matrix multiplications and convolutions in AI algorithms; AI VECTOR CORE accelerates vector operations through parallel processing; MIX AIC integrates different types of operators and optimizes for multiple tasks; AI CPU is optimized in hardware and instruction sets to better support AI algorithms. LocMoE+ selects fewer tokens, resulting in a 17$\times$ performance improvement in the FFN MatMul operator compared to the Baseline and a 2.6$\times$ improvement compared to LocMoE. This leads to an overall 2.8$\times$ reduction in the cumulative time consumption of the MatMul operator and a 2.6$\times$ decrease in Cube computing load. However, the proportions of TopK and IndexPutV2, related to rearrangement, show a slight increase.
Figure 4: The distribution of time consumption for operators.
Figure 5: The footprint recorded in one acquisition cycle with different schemes and cluster sizes.
We select a single iteration during the stable training period and describe the per-device memory usage (Allocated) using the first 100,000 samples from its memory monitoring, as shown in Figure 5. Overall, LocMoE+ achieves memory usage reduction of 4.57% to 16.27% compared to the Baseline and 2.86% to 10.5% compared to LocMoE. As cluster size increases, the proportion of computational overhead decreases, and the gap in memory usage narrows. Additionally, instantaneous memory peaks gradually disappear, and the fluctuation amplitude of short-term memory also diminishes.
4.3 The Performance of Downstream Tasks
To enhance the model’s conversational capabilities and adaptability to downstream task, we fine-tuned the pre-trained models. As shown in Figure 6, with sufficient supervised fine-tuning (SFT), LocMoE+ achieves an average improvement of approximately 20.1% in 16 sub-capabilities of Domain Task Capability, which is a portion of General and Domain-specific Assessment Dataset (GDAD), compared to the Baseline, and an increase of about 3.5% compared to LocMoE. The Rewriting and Summary capabilities show the highest improvement, with a 28.2% increase compared to the Baseline and a 6.7% increase compared to LocMoE. In the 13 tests of Domain Competency Exam, LocMoE+ demonstrates an average improvement of 16% relative to the Baseline and an average increase of approximately 4.8% compared to LocMoE. The IP Training in the digital communications domain shows the most significant improvement, with a 27.3% increase compared to the Baseline and a 3.0% increase compared to LocMoE. Among the 18 sub-capabilities of General Ability, LocMoE+ exhibits an improvement of about 13.9% relative to the Baseline and an average increase of 4.8% compared to LocMoE. The capability of Planning demonstrates the highest improvement, with a 26.8% increase compared to the Baseline and a 2.92% increase compared to LocMoE.
Figure 6: The performance on three categories of GDAD.
Table 1 presents the holistic evaluation results for multiple datasets, where GDAD-1 represents Domain Task Capability, and the other metrics follow accordingly. Notably, due to the 6:4 ratio of Chinese to English data in our incremental pre-training domain data and the 7:3 ratio in the fine-tuning data, LocMoE+ achieves an improvement of approximately 10.7% compared to the Baseline and 1.2% compared to LocMoE in the C-Eval [40] evaluation, despite the limited data available for training. During incremental training and fine-tuning, we incorporated substantial telecommunications domain knowledge, questions, and case studies. TeleQnA [41], the first benchmark dataset designed to evaluate the knowledge of LLMs in telecommunications, effectively measures the model’s capabilities in this domain. Consequently, LocMoE+ comprehensively surpasses both the Baseline and LocMoE on this specific dataset.
Table 1: Performance promotion obtained by LocMoE+ on different datasets.
GDAD
GDAD-1 GDAD-2 GDAD-3 Avg C-Eval TeleQnA
Baseline 47.8 43.0 65.4 52.8 38.5 62.1
LocMoE 55.5 47.6 71.1 59.0 42.6 67.6
LocMoE+ 57.4 49.9 74.5 61.5 43.1 68.8
5 Conclusion
In this paper, we propose a novel MoE structure incorporating token feature awareness mechanism to enhance the training efficiency and performance of LLMs. Building upon the feature extraction component and locality loss of LocMoE, LocMoE+ introduces a routing strategy that combines TCR and ECR, theoretically demonstrating improved training success rates. By defining the affinity between tokens and experts, expert capacity can be dynamically reduced and further minimized. Model training experiments are conducted on Ascend clusters, leveraging the communication optimization characteristics of CoC. LocMoE+ achieves performance improvements up to 46.6% (32N) compared to the Baseline and 13.3% (32N) compared to LocMoE, while reducing memory usage by up to 16.27% and 10.5%, respectively. To evaluate model performance, all pre-trained models are fine-tuned using a mixture of domain-specific and general data, and assessed with the open-source datasets C-Eval and TeleQnA, and closed domain benchmark GDAD. In downstream tasks, LocMoE+ outperforms the Baseline by 14.1%, 10.7%, and 9.7% on GDAD, C-Eval, and TeleQnA, respectively; it exceeds LocMoE by 4.1%, 1.2%, and 1.7% on these datasets, respectively. Both theoretical proofs and experimental results demonstrate the benefits of the LocMoE+ routing strategy for LLM training efficiency. Future work may explore methods to compress communication data to further reduce communication overhead.
Appendix A Missing Proofs
A.1 Auxiurary Results
Lemma A.1 (Theorem 4 in [42]).
Let ${\bm{X}}_{1},\dots,{\bm{X}}_{n}$ be $n$ independent random variables with
${\mathbb{P}}({\bm{X}}_{i}=1)=p_{i},{\mathbb{P}}({\bm{X}}_{i}=0)=1-p_{i}.$
We consider the sum ${\bm{X}}=\sum_{i=1}^{n}{\bm{X}}_{i}$, with expectation $\mathbb{E}({\bm{X}})=\sum_{i=1}^{n}p_{i}$. Then we have
$\mathrm{(Lower\ tail)}\quad{\mathbb{P}}({\bm{X}}\leq\mathbb{E}{\bm{X}}-\lambda)\leq e^{-\frac{\lambda^{2}}{2\mathbb{E}{\bm{X}}}},$ (9)
$\mathrm{(Upper\ tail)}\quad{\mathbb{P}}({\bm{X}}\geq\mathbb{E}{\bm{X}}+\lambda)\leq e^{-\frac{\lambda^{2}}{2(\mathbb{E}{\bm{X}}+\lambda/3)}}.$
A.2 Proof of Theorem 3.3
Proof.
1) For the TCR, denote $s_{i}=\left|\left\{t<k:{\bm{x}}_{t}\text{ sent to expert }i,{\bm{x}}_{k}={\bm{o}}_{i}\right\}\right|,\forall i\in[n]$ as the top class-irrelevant token number candidated to the $i$-th expert before the valid token. Then by Assumption 3.2, each class-irrelevant token uniformly gives to any expert, leading to $s_{i}|({\bm{x}}_{k}={\bm{o}}_{i})\sim{\mathcal{B}}(k-1,1/n)$ (Binomial distribution), that is,
${\mathbb{P}}(s_{i}=t|{\bm{x}}_{k}={\bm{o}}_{i})=\begin{pmatrix}k-1\\ t\end{pmatrix}\cdot\left(\frac{1}{n}\right)^{t}\left(1-\frac{1}{n}\right)^{k-1-t},\forall t\in[k-1].$
Then we could derive that
[…]
$=\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}{\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(C+\sum_{k=C+1}^{s}{\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})\right).$ (10)
Note that $\mathbb{E}s_{i}=(k-1)/n$. When $k\geq 2nC$, by lower tail bound in Lemma A.1, we get
${\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})\leq e^{-\frac{(k-1-n(C-1))^{2}}{2(k-1)n}}\leq e^{-\frac{k-1}{8n}}.$
Hence, we get the upper bound that
[…]
$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\sum_{k=2nC+1}^{s}{\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})\right)$
$\leq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\sum_{k=2nC}^{s-1}e^{-\frac{k}{8n}}\right)\leq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\frac{e^{-\frac{C}{4}}}{1-e^{-\frac{1}{8n}}}\right)$
$\stackrel{{\scriptstyle(i)}}{{\leq}}\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+(8n+1)e^{-\frac{C}{4}}\right)\leq\frac{10C\sum_{i=1}^{n}p_{i}}{s},$
where $(i)$ uses the inequality that $e^{-t}\leq 1/(1+t),\forall t\geq 0$.
Moreover, for $1+\frac{nC}{4}\leq k\leq 1+\frac{nC}{2}$, i.e., $2(k-1)\leq nC\leq 4(k-1)$, by upper tail bound in Lemma A.1, we get
${\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})=1-{\mathbb{P}}(s_{i}\geq C|{\bm{x}}_{k}={\bm{o}}_{i})\geq 1-e^{-\frac{3(nC-k+1)^{2}}{2n[2(k-1)+nC]}}\geq 1-e^{-\frac{k-1}{4n}}.$
Hence, we get the lower bound that
[…]
$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\sum_{k=\lceil 1+nC/4\rceil}^{\lfloor 1+nC/2\rfloor}{\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})\right)$
$\geq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\frac{nC}{4}-1-\sum_{k=\lceil 1+nC/4\rceil}^{\lfloor 1+nC/2\rfloor}e^{-\frac{k-1}{4n}}\right)\geq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\frac{nC}{4}-1-\frac{e^{-\frac{C}{16}}}{1-e^{-\frac{1}{4n}}}\right)$
$\stackrel{{\scriptstyle(i)}}{{\geq}}\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\frac{nC}{4}-2-(4n+1)e^{-\frac{C}{16}}\right)\geq\frac{C\sum_{i=1}^{n}p_{i}}{5s},$
where $(i)$ uses the inequality that $e^{-t}\leq 1/(1+t),\forall t\geq 0$, and the final inequality needs $C\geq 48$, which can be satisified in common experiments. Combining the upper and lower bounds, we obtain the desired result.
2) For the ECR, denote $s_{i}$ as the class-irrelevant token number with the score larger than ${\bm{o}}_{i}$ for $i$-th expert. By Assumption 3.2, we derive that $s_{i}\sim{\mathcal{B}}(s-1,q_{i}),\forall i\in[n]$.
[…]
$=\frac{1}{n}\sum_{i=1}^{n}{\mathbb{P}}(s_{i}\leq C-1,s_{i}\sim{\mathcal{B}}(s-1,q_{i}))$
If $C-1\leq(s-1)q_{i}/2$, by lower tail bound in Lemma A.1 with $\lambda=(s-1)q_{i}-(C-1)<\mathbb{E}s_{i}$, we obtain that
${\mathbb{P}}(s_{i}\leq C-1)\leq e^{-\frac{(s-1)q_{i}}{2}\left(1-\frac{C-1}{(s-1)q_{i}}\right)^{2}}\leq e^{-\frac{(s-1)q_{i}}{8}}.$
If $C\geq 2(s-1)q_{i}$, by upper tail bound in Lemma A.1 with $\lambda=C-(s-1)q_{i}>0$, we obtain that
${\mathbb{P}}(s_{i}\leq C-1)=1-{\mathbb{P}}(s_{i}\geq C)\geq 1-e^{-\frac{[C-(s-1)q_{i}]^{2}}{2(C+2(s-1)q_{i})/3}}\geq 1-e^{-\frac{3C}{16}}.$
Hence, we conclude Eq. (8). ∎
Appendix B Token Feature Distribution
We also validate the feature distribution before and after MoE training shown in Figure 7. We can see before training, all 8192 tokens in one training sample are nearly orthogonal with correlation coefficient near zero, which verifies the isotropy distribution assumption in the first bullet of Remark 3.4. After training, the token features are nearly aligned with correlation coefficien large than $0.8$. We can also observe that neighbouring tokens share similar features, and clear block feature behavior, meaning that the token features are relatively separated and the number of tokens in each cluster is bounded, which somehow matches the distribution assumption in the second bullet of Remark 3.4.
Figure 7: The correlation matrix of one training sample feature before (left) and after (right) training.
Appendix C Experimental Setup
C.1 Datasets for Training and Fine-Tuning
The dataset used in this paper is a self-constructed dataset that integrates knowledge from multiple domains, including wireless, data communication, and cloud-core technologies. It comprises Chinese, English, and bilingual corpora. The corpora are parsed from various internal technical documents, such as iCase, blogs, Wiki, and feature documents. Taking iCase as an example, iCase is a case record of problem localization and handling processes, containing code, instructions, and corresponding logs. In addition, the above-mentioned domain-specific knowledge corpora are mixed with general corpora in a ratio of 1:5. The general corpora are collected from hundreds of websites, including online novels, cooking guides, movie reviews, and more. After cleaning, deduplication, and review operations, the dataset is thoroughly shuffled. A total of 4.19 billion tokens is sampled as the experimental pre-training dataset. To evaluate downstream tasks, this paper also adopt hybrid sft data items to fine-tune the pre-trained model. The dataset comprises 762,321 general question-answer pairs and 11,048 domain-specific question-answer pairs, with a general-to-domain ratio of 68:1. The general characteristics encompass multi-tasking, mathematical ability, coding ability, logical reasoning, multi-turn dialogue, knowledge reasoning, language understanding, text generation, multi-tasking, FunctionCall, CoT, MRC summarization, refusal to answer, Chinese, and English. The domain-specific characteristics include domain knowledge understanding, RAG, FunctionCall, information extraction, multi-turn dialogue, reading comprehension, paraphrasing, and intent recognition.
C.2 Experimental Environment
The experiments are conducted on a cluster composed of Ascend 910B3 NPUs, divided into three groups: 32 NPUs (hereinafter referred to as 32N, and so on), 64N, and 256N. The 910B3 series NPU contains 20 AI cores with a main frequency of 1.8GHz and a theoretical computing power of 313T under fp16 precision. The physical High Bandwidth Memory (HBM) of the 910B3 NPU is 64G, with an HBM frequency of 1.6GHz and an HBM bandwidth of 1.6T. Every 8 NPUs are mounted on the same Atlas 800T A2 server, which internally adopts a fullmesh networking scheme, meaning that any two NPUs are interconnected. The version of the Ascend Hardware Development Kit (HDK) is 23.0.2.1, and the version of the Compute Architecture for Neural Networks (CANN) suite is 7.0.0, which is the commercial release version for Q4 2023. The models in this paper use ModelLink, an LLM training framework based on the Ascend architecture, and run in the torch_npu 5.0.0 environment.
C.3 Model Configuration
This paper adopts the Mixtral 8$\times$7B model with an MoE structure, released in December 2023, as the backbone and embeds LocMoE and LocMoE+. Mixtral has 46.7B parameters and uses the Group Query Attention (GQA) mechanism to compute attention. It contains 32 sparse expert blocks, with the MoE Feedforward layer comprising 8 experts, and each token selects the top 2 experts for processing. In the application scenarios of this paper, the corpora are generally long texts; thus, the sequence length is adjusted to 32768. For the three cluster scales of 32N, 64N, and 256N, the parallel strategies are set as follows: 32N - tensor parallel (TP=4) / pipeline parallel (PP=4) / data parallel (DP=2) / expert parallel (EP=2), 64N - TP=8 / PP=4 / DP=2 / EP=2, and 256N - TP=8 / PP=8 / DP=4 / EP=2. The batch size is set to 128.
C.4 Evaluation Metrics and Datasets
To evaluate model performance, this paper designs a comprehensive metric called the General and Domain-specific Assessment Dataset (GDAD), which consists of three evaluation systems: domain task capability, domain capability certification exam, and general capability. Among them, the domain task capability includes a total of 16 categories and 2,657 questions, such as domain logical reasoning; the domain capability certification exam includes a total of 13 categories and 13,968 questions, such as data communication; and the general capability includes a total of 18 categories and 1,435 questions, such as programming ability. The questions include objective and subjective questions in Chinese, English, and bilingual formats. For subjective questions, the cosine similarity between the model output and the standard answer is used as the score. In addition, this paper also employs C-Eval [40] and TeleQnA [41] to evaluate the model’s Chinese language capability.

### v2

@@
-LocMoE+: Enhanced Router with Token Feature Awareness for Efficient LLM Pre-Training
+Expert-Token Resonance: Redefining MoE Routing through Affinity-Driven Active Selection
 Abstract
-Mixture-of-Experts (MoE) architectures have recently gained increasing popularity within the domain of large language models (LLMs) due to their ability to significantly reduce training and inference overhead.
-However, MoE architectures face challenges, such as significant disparities in the number of tokens assigned to each expert and a tendency toward homogenization among experts, which adversely affects the model’s semantic generation capabilities.
-In this paper, we introduce LocMoE+, a refined version of the low-overhead LocMoE, incorporating the following enhancements: (1) Quantification and definition of the affinity between experts and tokens.
-(2) Implementation of a global-level adaptive routing strategy to rearrange tokens based on their affinity scores.
-(3) Reestimation of the lower bound for expert capacity, which has been shown to progressively decrease as the token feature distribution evolves.
-Experimental results demonstrate that, without compromising model convergence or efficacy, the number of tokens each expert processes can be reduced by over 60%.
-Combined with communication optimizations, this leads to an average improvement in training efficiency ranging from 5.4% to 46.6%.
-After fine-tuning, LocMoE+ exhibits a performance improvement of 9.7% to 14.1% across the GDAD, C-Eval, and TeleQnA datasets.
-1 Introduction
-Large language models (LLMs) have shown exceptional proficiency in understanding deep structures and complex semantic relationships within language [1].
-As these models scale up, their capabilities in language generation and logical comprehension are enhanced, but this comes at the cost of significant computational, communication, and storage demands [2].
-To scale models efficiently without disproportionately increasing computational costs, researchers have incorporated the Mixture-of-Experts (MoE) architecture into LLMs [3].
-The MoE framework integrates multiple experts within the model, each tasked with processing specific types of inputs [4].
-For a given input, only a subset of experts is activated, allowing for more efficient use of computational resources [5].
-Recently, several LLMs employing MoE structures, such as DeepSeek-MoE [6] and Mixtral [7], have demonstrated outstanding performance on various leaderboards.
-Despite the efficiency benefits of MoE in scaling model sizes, it introduces several new challenges and drawbacks [8].
-For instance, an imbalanced "winner-takes-all" phenomenon among experts can detrimentally affect overall model performance [9], and the All-to-All communication pattern can become a bottleneck [10].
-The model’s convergence and the experts’ generalization capabilities are heavily dependent on the design of the routing strategy.
-Recent research has addressed these challenges from multiple angles, such as optimizing routing strategies and enhancing communication efficiency.
-In terms of routing strategies, DeepSeek-MoE [6] introduced the concepts of vertical-class experts, created through fine-grained partitioning, and shared experts, which are always activated.
-Google Brain proposed the Expert Choice (EC) routing algorithm [11], assigning experts with predetermined buffer capacities to the Top-k tokens to ensure load balance.
-For communication efficiency, FastMoE [12] leverages Megatron-LM [13], a distributed training framework by NVIDIA, utilizing CUDA’s multi-stream mechanism to maximize the overlap between communication and computation.
-Tutel [14] introduced flexible All-to-All communication to alleviate throughput bottlenecks caused by smaller matrix dimensions and proposed 2DH All-to-All to align non-contiguous memory spaces and merge scattered communications.
-The uneven expert load in classical MoE [15] often necessitates token dropping or padding.
-Megablocks [16] employs block-sparse operations to overcome the limitations of dynamic routing, improving the balance between model quality and efficiency.
-LocMoE [17] aims to enhance the interpretability of routing mechanisms while reducing computational and communication overhead.
-LocMoE leverages orthogonal routing weights to prevent token homogenization across different expert networks and introduces the Grouped Average Pooling (GrAP) layer [18] for token feature extraction.
-Under these conditions, LocMoE uses theoretical proofs to calculate the lower bound of expert capacity.
-Experimental results demonstrate that employing this lower bound does not degrade, but slightly enhances, performance on downstream tasks.
-Additionally, LocMoE introduces a loss regularization term that incorporates data locality and load balancing [19], leading to a 12% to 22% improvement in training efficiency.
-Building on previous work, this paper proposes LocMoE+, which aims to further enhance training efficiency and model performance.
-LocMoE+ adopts GrAP and the locality loss regularization terms from LocMoE, and continues to analyze the impact of token features on expert selection.
+Mixture-of-Experts (MoE) architectures have emerged as a paradigm-shifting approach for large language models (LLMs), offering unprecedented computational efficiency.
+However, these architectures grapple with challenges of token distribution imbalance and expert homogenization, impeding optimal semantic generalization.
+We introduce a novel framework that redefines MoE routing through affinity-driven active selection.
+The innovations for the framework encompass: (1) A rigorous formulation of expert-token affinity metrics.
+(2) An adaptive bidirectional selection mechanism leveraging resonance between experts and tokens.
+(3) Theoretical derivation and experimental evidence of reduced expert capacity bounds under dynamic token distribution evolution.
+It is also integrated with orthogonal feature extraction module and an optimized loss function for expert localization.
+Our theoretical analysis demonstrates that this approach mitigates expert homogenization while enabling substantial capacity boundary reduction.
+Experimental validation corroborates these findings: it achieves a 40% reduction in token processed by each expert without compromising model convergence or efficacy.
+When coupled with communication optimizations, the training efficiency improvements of 5.4% to 46.6% can be observed.
+After supervised fine-tuning, it exhibits performance gains of 9.7% to 14.1% across GDAD, C-Eval, and TeleQnA benchmarks.
+Introduction
+Large language models (LLMs) have shown exceptional proficiency in understanding deep structures and complex semantic relationships within language (Zhao et al.
+2023).
+As these models scale up, their capabilities in language generation and logical comprehension are enhanced, but this comes at the cost of significant computational, communication, and storage demands (Jiang et al.
+2024b).
+To scale models efficiently without disproportionately increasing computational costs, researchers have incorporated the Mixture-of-Experts (MoE) architecture into LLMs (Lepikhin et al.
+2020).
+The MoE framework integrates multiple experts within the model, each tasked with processing specific types of inputs (Fedus, Zoph, and Shazeer 2022).
+For a given input, only a subset of experts is activated, allowing for more efficient use of computational resources (Du et al.
+2022).
+Recently, several LLMs employing MoE structures, such as DeepSeek-MoE (Dai et al.
+2024) and Mixtral (Jiang et al.
+2024a), have demonstrated outstanding performance on various leaderboards.
+Despite the efficiency benefits of MoE in scaling model sizes, it introduces several new challenges and drawbacks (Shazeer et al.
+2017).
+The conventional MoE model’s convergence and the experts’ generalization capabilities are heavily dependent on the design of the routing strategy, which easily leads to an imbalanced ”winner-takes-all” phenomenon among experts.
+The imbalance between excessively ”developed” experts and those lacking adequate training may compromise or even nullify the intended functionality of routing strategies.
+Recent studies address these challenges from multiple perspectives (Li et al.
+2023).
+DeepSeek-MoE (Dai et al.
+2024) introduces the concept of the vertical-class expert, which is created through fine-grained partitioning, and the shared experts are always activated.
+This approach minimizes the constraints on expert specialization imposed by knowledge hybridity and knowledge redundancy.
+StableMoE (Dai et al.
+2022) proposes a two-stage training approach to address the issue of routing fluctuation.
+This method involves training the routing network independently from the backbone model and utilizing a frozen, distilled routing mechanism to allocate tokens.
+The characteristics of classical gated routing lead to experts being unable to learn features mastered by other experts.
+To address this, MoDE (Xie et al.
+2024) proposes moderate distillation between experts to mitigate the generalization problems caused by narrow learning paths.
+DYNMoE (Guo et al.
+2024) introduces a unique gated routing mechanism capable of adaptively determining the number of activated experts through trainable expert thresholds, even allowing for the addition or removal of experts.
+In addition to the classical token choice scenario, previous researches also propose work utilizing expert choice (EC).
+Google Brain introduces the EC routing algorithm (Zhou et al.
+2022), which assigns experts with predetermined buffer capacities to the Top-k tokens to ensure load balance.
+The Brainformer (Zhou et al.
+2023) also adopts this routing strategy, constructing a trainable gating matrix to project the input feature space onto scores corresponding to each expert.
+Then, each token is routed to the Top-k experts.
+This strategy is proven highly effective in achieving expert load balancing and enhancing expert learning outcomes.
+The design of routing strategy is crucial to the MoE structure, while not all tokens may be suitable for training (Riquelme et al.
+2021).
+In addition to data preprocessing techniques such as dataset cleaning and deduplication, previous studies have also considered how to discard certain tokens within the model.
+Early work introduced the concept of expert capacity (Lepikhin et al.
+), which refers to the maximum number of tokens each expert can process at once.
+Tokens exceeding this capacity are discarded.
+Expert capacity helps to ensure load balance among experts while facilitating All-to-All communication implementation.
+However, in situations where it is uncertain whether a token contributes to training, there is a risk of discarding class-discriminative samples, potentially compromising the model’s training outcomes.
+XMoE (Yang et al.
+2024) achieves more precise router by implementing a threshold-based approach.
+If a token reaches the specified threshold, it is processed exclusively by a single expert while being discarded by other experts within the Top-k selection.
+This method allows for more nuanced token selection and processing.
+LocMoE (Li et al.
+2024) leverages orthogonal routing weights to prevent token homogenization across different expert networks and introduces the Grouped Average Pooling (GrAP) layer (Wang, Zhang, and Du 2023) for token feature extraction.
+Under these conditions, LocMoE also provides the theoretical proof for the lower bound of expert capacity.
+In this paper, we propose expert-token resonance, a mechanism consisting of an expert-token bidirectional selection router and the adaptive expert capacity strategy.
 The primary contributions of this paper are as follows:
   1. 1.
-    The affinity score.
+    A rigorous formulation of expert-token affinity metrics.
 Given that expert selection is related to the angle between tokens and gating weights, the cosine similarity between tokens and gating weights is used to define the affinity score between tokens and experts.
-This score is utilized for token rearrangement and as a threshold for the lower bound of expert capacity.
+We provide theoretical analysis of utilizing the affinity score for better guiding expert choice routers to focus on different segmentations of tokens, thus mitigating the homogenization tendency.
+Further, we incorporate affinity scores into the design of our bidirectional selection router.
   2. 2.
-    The hybrid token rearrangement scheme.
-By integrating the concepts of Token Choice Router (TCR) and Expert Choice Router (ECR), the training success rate is enhanced while considering expert capacity constraints.
+    Expert-token bidirectional selection.
+By integrating the concepts of expert choice router (ECR) and token choice router (TCR), we propose the adaptive bidirectional selection mechanism.
+Contrast to conventional router, the bidirectional selection router allows MoE to enhance the training success rate while considering expert capacity constraints.
 Its effectiveness has been theoretically validated.
   3. 3.
-    The adaptive expert capacity.
-Setting a holistic affinity threshold allows the lower bound of expert capacity to be significantly reduced.
-As training iterations increase, the information density of token features grows, causing the expert capacity to initially decrease and then stabilize, thereby enhancing training efficiency.
-LocMoE+ uses the state-of-the-art MoE model Mixtral 8$\times$7B as the backbone, and utlizes AscendSpeed and ModelLink libraries for pre-training on Ascend NPU clusters.
-Experiments conducted on clusters with 32, 64, and 256 NPUs show that LocMoE+ improves training efficiency by 5.4% to 46.6% compared to the baseline, and by 2.9% to 13.3% compared to LocMoE.
+    The adaptive expert capacity bound.
+Setting an adaptive affinity threshold allows the lower bound of expert capacity to be significantly reduced.
+As training iterations increase, the information density of token features grows, causing the expert capacity to initially decrease and then stabilize.
+Ultimately, the training efficiency of MoE can be greatly enhanced.
+Expert-token resonance mechanism adopts the state-of-the-art MoE model Mixtral 8$\times$7B as the backbone, and utilizes AscendSpeed and ModelLink libraries for pre-training on Ascend NPU clusters.
+Experiments conducted on clusters with 32, 64, and 256 NPUs indicate that our approach improves training efficiency by 5.4% to 46.6% compared to the baseline, and by 2.9% to 13.3% compared to LocMoE.
 Model performance is enhanced by 9.7% to 14.1% compared to the baseline, and by 1.7% to 4.1% compared to LocMoE.
-The rest of this paper is structured as follows: Section 2 introduces related work on LLMs, MoE, and the Ascend architecture.
-Section 3 reviews LocMoE and presents the methods proposed in this paper, along with theoretical evidence.
-Section 4 analyzes the experimental results of LocMoE+ regarding training efficiency and model performance.
+The rest of this paper is structured as follows: Section Related Work introduces related work on LLMs, MoE, and the Ascend architecture.
+Section Method reviews LocMoE and presents the methods proposed in this paper, along with theoretical evidence.
+Section Experiments analyzes the experimental results of our approach regarding training efficiency and model performance.
 The final section summarizes the content of this paper and offers an outlook on future improvements.
-2 Related Work
+Related Work
 LLMs.
-BERT [20] and GPT [21] in 2018 marked the emergence of the LLM era.
-GPT-3 [22], with 175 billion parameters, demonstrated impressive capabilities in language understanding and generation in 2020.
-By 2023-2024, LLMs expanded to hundreds of billions to trillions of parameters: GPT-4 [23] achieved breakthroughs in zero-shot and few-shot learning; PaLM-2 [24] enhanced knowledge acquisition and reasoning; Chinchilla [25] optimized computational resources and model scale; and Claude-v1 [26] displayed a safer, human-aligned LLM.
+BERT (Kenton and Toutanova 2019) and GPT (Radford et al.
+) in 2018 marked the emergence of the LLM era.
+GPT-3 (Floridi and Chiriatti 2020), with 175 billion parameters, demonstrated impressive capabilities in language understanding and generation in 2020.
+By 2023-2024, LLMs expanded to hundreds of billions to trillions of parameters: GPT-4 (Achiam et al.
+2023) achieved breakthroughs in zero-shot and few-shot learning; PaLM-2 (Anil et al.
+2023) enhanced knowledge acquisition and reasoning; Chinchilla (Hoffmann et al.
+2022) optimized computational resources and model scale; and Claude-v1 (Wei, Haghtalab, and Steinhardt 2024) displayed a safer, human-aligned LLM.
 These LLMs have made significant advancements across domains, paving the way towards artificial general intelligence.
@@
 To alleviate this issue, researchers have introduced the MoE architecture, which incorporates multiple expert modules and dynamically selects a subset of experts to process inputs, significantly reducing computational overhead while maintaining performance.
-GShard [3] and Switch Transformer [4] employ Top-2 and Top-1 routing strategies, respectively, to mitigate the expert load imbalance problem.
-Hash Layer [27] adopt hashing techniques to simplify the gating mechanism and improve model performance.
-METRO [28] introduces a metric learning-based routing mechanism that guides expert selection by learning task-relevant metric functions, enhancing the model’s generalization ability.
-DSelect-k [29] proposes a dynamic selection algorithm based on Markov chains, adaptively choosing the number of experts to maintain performance while reducing training costs.
-Furthermore, DeepSpeed-MoE [30] optimizes expert parallelism and pipeline parallelism in MoE, further boosting training efficiency.
+GShard (Lepikhin et al.
+2020) and Switch Transformer (Fedus, Zoph, and Shazeer 2022) employ Top-2 and Top-1 routing strategies, respectively, to mitigate the expert load imbalance problem.
+Hash Layer (Roller et al.
+2021) adopt hashing techniques to simplify the gating mechanism and improve model performance.
+METRO (Bajaj et al.
+2022) introduces a metric learning-based routing mechanism that guides expert selection by learning task-relevant metric functions, enhancing the model’s generalization ability.
+DSelect-k (Hazimeh et al.
+2021) proposes a dynamic selection algorithm based on Markov chains, adaptively choosing the number of experts to maintain performance while reducing training costs.
+Furthermore, DeepSpeed-MoE (Rajbhandari et al.
+2022) optimizes expert parallelism and pipeline parallelism in MoE, further boosting training efficiency.
 These works lay the foundation for the subsequent application of MoE in LLMs.
 Ascend Architecture.
-The AI Core in the Ascend 910B NPU adopts the DaVinci architecture [31], which is responsible for executing computation-intensive operators related to scalars, vectors, and 3D cubes, and is widely used in LLM training.
-Huawei’s proprietary cache coherence protocol interconnect, High Confidence Computing Systems (HCCS) [32], enables high-performance inter-device data communication in multi-card scenarios.
-The Huawei Collective Communication Library (HCCL) [33], designed for the Ascend processor, offers high-performance primitives for single-node multi-card and multi-node multi-card collective communication.
+The AI Core in the Ascend 910B NPU adopts the DaVinci architecture (Tang and Wang 2023), which is responsible for executing computation-intensive operators related to scalars, vectors, and 3D cubes, and is widely used in LLM training.
+Huawei’s proprietary cache coherence protocol interconnect, High Confidence Computing Systems (HCCS) (Xia et al.
+2021), enables high-performance inter-device data communication in multi-card scenarios.
+The Huawei Collective Communication Library (HCCL) (Cao et al.
+2024), designed for the Ascend processor, offers high-performance primitives for single-node multi-card and multi-node multi-card collective communication.
 It implements collective communication functions on PCIe, HCCS, and RoCE high-speed links, facilitating distributed training.
-PyTorch [34], a widely adopted deep learning framework, is compatible with the Ascend NPU through the Ascend extension for PyTorch [35].
+PyTorch (Paszke et al.
+2019), a widely adopted deep learning framework, is compatible with the Ascend NPU through the Ascend extension for PyTorch (Zhu et al.
+2023).
 AscendSpeed, an acceleration library tailored for Ascend devices, focuses on parallel strategy partitioning for LLMs and aggregates results through collective communication.
 ModelLink, a training framework designed for LLMs, fully considers the characteristics of the Ascend NPU and implements features such as memory optimization, multi-dimensional parallelism, and adaptive data loading, further enhancing training performance.
-3 Method
-Notation.
-[…]
-3.1 Background
-Main Principle.
+Method
+In this section, we present the definition of affinity score with a rigorous formulation of expert-token affinity metrics and the implementation of our affinity-driven active selection routing.
+To reduce the lower bound of expert capacity, we introduce a training strategy for dynamic token distribution evolution under the guiding principle of affinity score.
+Model Architecture.
+Backbone.
 The MoE architecture, based on the Transformer framework, efficiently scales up model size with low computational overhead, benefiting from two primary structures: a sparse gating network for routing tokens and expert networks for processing specific token categories.
-The sparse gating network determines the expert to which a token will be sent.
 We consider the supervised classification for brevity where the training samples are $\{(\bm{x}^{(i)},y_{i})\}_{i=1}^{N}\sim\mathcal{D}$.
-Each training sample ${\bm{x}}^{\top}=({\bm{x}}_{1}^{\top},\dots,{\bm{x}}_{s}^{\top})\in{\mathbb{R}}^{sd}$ has $s$ tokens with token feature ${\bm{x}}_{i}\in{\mathbb{R}}^{d},\forall i\in[s]$, and label $y\in{\mathbb{N}}^{+}$.
+Each training sample $\bm{x}^{\top}=(\bm{x}_{1}^{\top},\dots,\bm{x}_{s}^{\top})\in\mathcal{R}^{sd}$ has $s$ tokens with token feature $\bm{x}_{i}\in\mathcal{R}^{d},\forall i\in[s]$, and label $y\in\mathcal{N}^{+}$.
 The objective is to learn the map of $\bm{x}$ to the corresponding $y$.
@@
 $\text{MoE}(\bm{x})=\sum_{t=1}^{s}\sum_{i=1}^{n}G_{i}(\bm{x}_{t})\cdot E_{i}(\bm{x}_{t}),$ (1)
-where $n$ is the number of experts, $G({\bm{x}}_{t})\colon{\mathbb{R}}^{d}\to{\mathbb{R}}^{n}$ is the gating (or router) weights of experts which maps the tokens of ${\bm{x}}_{t}$ into the coresponding experts with weights, e.g., $G_{i}({\bm{x}})=\text{Softmax}({\bm{W}}{\bm{x}}+\bm{\epsilon})$ where the softmax is applied to each row, and $E_{i}({\bm{x}}_{t})\colon{\mathbb{R}}^{d}\to{\mathbb{R}}$ is the $i$-th expert network, e.g., simple MLP models, see [36] for current different router methods.
+where $n$ is the number of experts, $G(\bm{x}_{t})\colon\mathcal{R}^{d}\to\mathcal{R}^{n}$ is the gating weight vector of experts which maps the tokens of $\bm{x}_{t}$ into the coresponding experts with weights, e.g., $G_{i}(\bm{x})=\text{Softmax}(\bm{W}\bm{x}+\bm{\epsilon})$ where the softmax is applied to each row, and $E_{i}(\bm{x}_{t})\colon\mathcal{R}^{d}\to\mathcal{R}$ is the $i$-th expert network, see (Liu et al.
+2024) for current different router methods.
 Generally, $n\ll s$, which saves much computation compared to the dense structure.
+Expert-Token Affinity.
+$\bm{W}_{\text{aff}}$ denotes the expert-token affinity matrix. With GrAP as the layer of feature extraction, the formulation of $\bm{W}_{\text{aff}}$ is as followed:
+$\bm{W}_{\text{aff}}=\begin{pmatrix}\bm{w}_{1}&0&\cdots&0\\ 0&\bm{w}_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\bm{w}_{n}\\ \end{pmatrix}$ (2)
+$\bm{w}_{i}=\frac{n}{d}\cdot\mathbf{1}\left\{\frac{i\cdot d}{n}\leq j<\left(i+1\right)\frac{d}{n}\right\}\quad 0\leq j<d$ (3)
+The expert-token affinity matrix is employed as the gating weight to calculate the affinity score between each expert and token.
+We define the affinity score of $t$-th token and $i$-th expert as the cosine similarity between vectors $\bm{x}_{t}$ and $\bm{w}_{i}$:
+$\delta_{ti}=\cos\left(\bm{x}_{t},\bm{w}_{i}\right):=\bm{x}_{t}^{\top}\bm{w}_{i}/(\|\bm{x}_{t}\|\cdot\|\bm{w}_{i}\|)$ (4)
+The affinity score intuitively reflect how closely the two inputs are associated.
+From a perspective of semantic, the affinity scores derived from affinity metrics consisting of orthogonal vectors represent the degree of association between each token and various experts, as shown in Figure 1.
+Therefore, we leverage the affinity score as the principle of our affinity-driven active selection routing mechanism.
+Figure 1: The illustration of affinity score.
+Figure 2: The architecture of the gate network along with the hybrid TCR + ECR router.
 Routing Strategy.
-There are two main router methods: hard router [36] which uses hard assignments between tokens and experts with sparse dispatch matrix, and soft router [37] which uses weighted combinations of tokens.
-In this paper, we still focus on the first one, since the computation cost is much smaller than the soft ones.
-The hard router also have two main choices: 1) Token Choice Router (TCR) [38, 11], which lets each token choose its top-scored experts; Taking Top1 routing as an example, it typically selects a linear layer to extract feature representations for different experts and then obtains the expert index with the highest routing probability through the gating function.
-2) Expert Choice Router (ECR), which lets each expert choose its top-scored tokens.
-Moreover, to balance the budget of differnet experts, we will pre-define the expert capacity $C$ of all experts in practice, namely, the maximum tokens sent to each expert.
-Chocie of score.
-In this paper, we define the score of $t$-th token and $i$-th expert as $\delta_{ti}=\cos\left({\bm{x}}_{t},{\bm{w}}_{i}\right):={\bm{x}}_{t}^{\top}{\bm{w}}_{i}/(\|{\bm{x}}_{t}\|\cdot\|{\bm{w}}_{i}\|)$, i.e., the cosine simlarity between vectors ${\bm{x}}_{t}$ and ${\bm{w}}_{i}$.
-The score is to decide the gating weight in router.
-For example, we can route $t$-th token to the experts of the $-\ell$ indices in $(\delta_{t1},\dots,\delta_{tn})$.
-Chocie of capacity.
-In previous works, expert capacity $C$ is calculated using a capacity factor, ensuring that the size of each expert’s token buffer remains consistent.
-The choice of capacity factor is currently based mainly on empirical knowledge; increasing the capacity factor allows for handling more tokens but reduces training efficiency.
-TCR expression. First, given $\ell$, we decide the candidate expert set for each token by the sort of scores:
-$\left(\tilde{I}_{t1},\dots,\tilde{I}_{t\ell}\right)=\text{Top-}\ell\left(\left\{\delta_{t1},\dots,\delta_{tn}\right\}\right),\tilde{I}_{tk}\in[n],\forall t\in[s],k\in[\ell].$ (2)
-Second, we choose the first-$C$ tokens (or all if less than $C$) to each expert due to the capacity limit:
-$\left(I_{1i},\dots,I_{Ci}\right)=\text{Bottom-}C\left(\left\{t\in[s]:\exists j\in[\ell],\tilde{I}_{tj}=i\right\}\right),I_{ki}\in[s]\cup\text{None},\forall i\in[n],k\in[C].$ (3)
-In practice, we adopt $\ell=1$ to further save the computation cost since $n\ll s$.
-ECR expression. We directly choose the tokens in the top-$C$ scores of each expert:
-$\left(I_{1i},\dots,I_{Ci}\right)=\text{Top-}C\left(\left\{\delta_{1i},\dots,\delta_{si}\right\}\right),I_{ki}\in[s],\forall i\in[n],k\in[C].$ (4)
-Intuitively, TCR can result in underused experts with fewer tokens than their allowed capacity, while ECR can result in underused tokens where no expert receives such token.
-Auxiliary Loss.
+We consider our affinity-driven active selection routing as a hybrid of TCR (Clark et al.
+2022; Zhou et al.
+2022) and ECR.
+As the name suggested, TCR lets each token choose its top-scored experts, and ECR lets each expert choose its top-scored tokens.
+Specifically, we use the result of the expert-token affinity metrics as the affinity score between tokens and experts.
+In conventional TCR routing strategy, the tokens are simply route to their Top-1 expert.
+In our hybird TCR+ECR routing strategy, experts also select tokens for processing from assigned tokens according to affinity scores:
+$\left(\tilde{E}_{t1},\dots,\tilde{E}_{t\ell}\right)=\text{Top-}\ell\left(\left\{\delta_{t1},\dots,\delta_{tn}\right\}\right),$ (5)
+$~{}~{}\tilde{I}_{tk}\in[n],\forall t\in[s],k\in[\ell].$
+and then the expert to choose its $-\ell$ tokens where $\ell$ is determined by a threshold of the sum of affinity scores:
+$\left(I_{1i},\dots,I_{Ci}\right)=\text{Bottom-}C\left(\left\{t\in[s]:\exists j\in[\ell],\tilde{I}_{tj}=i\right\}\right),$ (6)
+$~{}~{}I_{ki}\in[s]\cup\text{None},\forall i\in[n],k\in[C].$
+Such bidirectional selection mechanism motivates each expert to receive a certain number of tokens with the highest affinity score to itself, thereby achieving a resonance effect.
+The resonance effect can help mitigate the homogenization in MoE.
+Locality Loss.
 Feed-forward network (FFN) layers are commonly employed in expert networks, allowing each expert to learn independently as a separate neural network, thus preventing interference between samples.
 This mechanism leads to a severe load imbalance, as experts frequently selected in the early stages are more likely to be chosen in later stages.
-To mitigate this skewness in token allocation, the auxiliary loss [8] has been proposed, defined as $L_{\mathrm{aux}}=\alpha{{n}{\sum_{i=1}^{n}f_{i}{P_{i}}}}$, where $f_{i}=\frac{1}{T}\sum_{{\bm{x}}\in{\beta}}\mathds{1}{\mathop{\arg\max}p({\bm{x}})=i}$, and $P_{i}=\frac{1}{T}\sum{{\bm{x}}\in{\beta}}p_{i}({\bm{x}})$.
-Here, $\beta$ represents the batch containing $T$ tokens, and $p_{i}$ denotes the probability of a token being routed to expert $i$.
-3.2 LocMoE
-GrAP.
-[…]
-The Lower Bound of the Expert Capacity.
+To mitigate this skewness in token allocation, the auxiliary loss (Shazeer et al.
+2017) has been proposed.
+Building upon the auxiliary loss, LocMoE introduces a loss bias term based on data locality, represented as $L_{\mathrm{loc}}=\mu\mathrm{KL}(D_{\mathrm{c}}||D_{\mathrm{l}})=-\mu\int{D_{\mathrm{c}}(x)\ln[\frac{D_{\mathrm{l}}(x)}{D_{\mathrm{c}}(x)}]\mathrm{d}x}$, i.e., the Kullback-Leibler (KL) divergence of the current distribution $D_{\mathrm{c}}(x)$ and the fully localized distribution $D_{\mathrm{l}}(x)$.
+This loss term serves as a soft constraint, encouraging tokens to be sent to experts residing on the same node, thereby mitigating the substantial overhead incurred by partial inter-node communication.
+Training Strategy
+The Dynamic Token Distribution Evolution.
 Under the premises of orthogonal gating weights and a data distribution approaching uniformity, LocMoE demonstrates that the expert capacity is closely related to the angle between the gating weights and tokens.
 For large scale of the activation, the lower bound of expert capacity is proven to exist and is represented as $C_{\mathrm{min}}=\frac{1}{n}\exp\{d\delta_{\mathrm{max}}^{2}/(2-\delta_{\mathrm{max}}^{2})\}$.
-Locality Loss.
-Building upon the auxiliary loss, LocMoE introduces a loss bias term based on data locality, represented as $L_{\mathrm{loc}}=\mu\mathrm{KL}(D_{\mathrm{c}}||D_{\mathrm{l}})=-\mu\int{D_{\mathrm{c}}(x)\ln[\frac{D_{\mathrm{l}}(x)}{D_{\mathrm{c}}(x)}]\mathrm{d}x}$, i.e., the Kullback-Leibler (KL) divergence of the current distribution $D_{\mathrm{c}}(x)$ and the fully localized distribution $D_{\mathrm{l}}(x)$.
-This loss term serves as a soft constraint, encouraging tokens to be sent to experts residing on the same node, thereby mitigating the substantial overhead incurred by partial inter-node communication.
-3.3 Token Awareness Router for LocMoE
-3.3.1 Scheme Sketch
-Figure 1 illustrates the model constructed in this paper, with a particular focus on the proposed token awareness router.
-The model framework is GPT-like, comprising only decoders, and incorporates LocMoE+ in place of the gated routing of sparse MoE blocks.
-To enhance interpretability and reduce computational overhead, the original Top-2 routing is simplified to the sparsest Top-1 routing.
-Consequently, the weights of GrAP, consisting of orthogonal vectors, emphasize different aspects of token features, which are evenly divided and non-overlapping.
-This routing mechanism can be viewed as a multi-classification of these feature fragments.
-TCR, introduced in the background, is exemplified in the figure.
+The hybrid TCR+ECR bidirectional selection routing, introduced in the model structure, is exemplified in the figure.
 If the feature fragment corresponding to the $k$-th dimension of the gating weight for a particular token is more prominent, then that token will be routed to the $k$-th expert.
@@
 Subsequent theoretical analysis will demonstrate the effectiveness of this hybrid routing scheme.
-Figure 1: The architecture of the model along with the hybrid TCR+ECR router.
-3.3.2 Theoretical Explanation
-To explain the motivation of our method, we show some theoretical insights in this section.
-Our theoretical analyis is bulit on [39], where they assume the training data has unique class-discriminative patterns with exectly the same number as experts, and other class-irrelevant patterns.
-Chowdhury et al.
-[39] demonstrated that the classical MoE structure go through two phases:
-Phase 1: Router training ([39, Lemma 4.1, Assumption 4.4]), which makes class-discriminative patterns all to the corresponding expert.
+Figure 3: (a) The average composition of computation, communication, overlap, and idle with different schemes and cluster sizes. (b) The perplexity during training iterations with different schemes.
+Theoretical Explanation
+To explain the motivation of our method, we show some theoretical insights in this section. Our theoretical analyis is bulit on Chowdhury et al. (2023), where they make the following data assumption:
+Assumption 1 (data assumption).
+Each input $\bm{x}\in\mathcal{R}^{sd}$ with $s$ tokens is comprised of one class-discriminative pattern $\bm{o}_{1},\dots,\bm{o}_{n}\in\mathcal{R}^{d}$, with each decides the label in $[n]$, and $s-1$ class-irrelevant patterns $\bm{r}\sim\mathcal{N}$ for certain distribution $\mathcal{N}$.
+For example, $\bm{x}=(\bm{r}_{1},\bm{r}_{2},\bm{o}_{1},\bm{r}_{3},\dots,\bm{r}_{s-1})$ has label $1$, where $\bm{r}_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N},\forall i\in[s-1]$.
+Based on Assumption 1, Chowdhury et al. (2023) demonstrated that the training of MoE go through two phases:
+Phase 1: Router training (Chowdhury et al.
+2023, Lemma 4.1 and Assumption 4.4), which makes class-discriminative patterns all to the corresponding expert.
 This process ensures that each expert only receives the class-discriminative tokens related to the specific class.
-Phase 2: Expert training ([39, Theorem 4.2, Theorme 4.5]), which makes each expert learn to predict the label based on its class-discriminative inputs from Phase 1.
+Phase 2: Expert training (Chowdhury et al.
+2023, Theorem 4.2 and Theorme 4.5), which makes each expert learn to predict the label based on its class-discriminative inputs from Phase 1.
 This process is designed to establish each expert’s ability to handle and solve problems.
 Hence, the traning of an input in the current step is valid if the class-discriminative patterns is correctly dispatched.
-We compute training success rate of input (i.e., the probability of the tokens in an input is correctly dispatched to experts) to measure the difference between TCR and ECR,
-We follow the setup in [39] that the input ${\bm{x}}\in{\mathbb{R}}^{sd}$ with $s$ tokens is comprised of one class-discriminative pattern ${\bm{o}}_{1},\dots,{\bm{o}}_{n}\in{\mathbb{R}}^{d}$, with each decides the label $y\in[n]$, and $s-1$ class-irrelevant patterns ${\bm{r}}\sim{\mathcal{N}}$ (the distribution will be clarified below).
-To show the quantitative comparison of TCR and ECR in training success rate, we need following asssumptions of token patterns.
-Assumption 3.1 (class-discriminative).
+To quantitatively measure the difference between TCR and ECR, we define training success rate of input motivated by the training process of MoE.
+Definition 2 (training success rate).
+We say the input $\bm{x}\in\mathcal{R}^{sd}$ with $s$ tokens succeed in training if the class-discriminative pattern in $\bm{x}$, e.g., $\bm{o}_{i}$ is correctly dispatched to $i$-th expert.
+We further define training success rate as the probability that the input succeed in training.
+Furthermore, to show the quantitative comparison of TCR and ECR in training success rate, we need following asssumptions and notations of token patterns.
+Assumption 3 (class-discriminative).
 We assume the location and feature of class-discriminative pattern is uniformly distribute in $[s]$ and $[n]$, i.e.,
@@
 We also assume that $\forall i\in[n],\bm{o}_{i}$ should be sent to the $i$-th expert, and define the true positive probability in token choice setting is no worse than the uniform dispatch as below
-${\mathbb{P}}({\bm{o}}_{i}\text{ sent to expert }i)=p_{i}\geq 1/n,\forall i\in[n].$
-Assumption 3.2 (class-irrelevant).
+$\mathcal{P}(\delta_{\bm{o}_{i},i}\geq\delta_{\bm{x}_{j},i},\forall j\in[s])=p_{i}\geq 1/n,\forall i\in[n].$
+Assumption 4 (class-irrelevant).
 The distribution of class-irrelevant patterns is isotropy, i.e.,
-${\mathbb{P}}({\bm{r}}\sim{\mathcal{N}},{\bm{r}}\text{ sent to expert }i)=1/n,\forall i\in[n].$ (5)
+$\mathcal{P}(\bm{r}\sim\mathcal{N},\delta_{\bm{r},i}\geq\delta_{\bm{x}_{j},i},\forall j\in[s])=1/n,\forall i\in[n].$ (7)
 And we define the false positive probability in expert choice setting as
-${\mathbb{P}}({\bm{r}}\sim{\mathcal{N}},\text{ the score of }{\bm{r}}\text{ is larger than }{\bm{o}}_{i}\text{ for expert }i)=q_{i},\forall i\in[n],$ (6)
+$\mathcal{P}(\bm{r}\sim\mathcal{N},\delta_{\bm{r},i}\geq\delta_{\bm{o}_{i},i})=q_{i},\forall i\in[n],$ (8)
 which measures the possibility that expert $i$ chooses the wrong token $\bm{r}$ instead of the correct token $\bm{o}_{i}$.
-Assumption 3.1 assumes the valid token is uniformly distributed in training samples due to the massive amounts of data nowadays.
-Assumption 3.2 assumes the invalid tokens can be uniformly dispatched to experts since the invalid tokens do not provide supervised signal to router and experts in training.
+Assumption 3 assumes the valid token is uniformly distributed in training samples due to the massive amounts of data nowadays.
+Assumption 4 assumes the invalid tokens can be uniformly dispatched to experts since the invalid tokens do not provide supervised signal to router and experts in training.
 We consider such uniform settings are common assumptions in theoretical analyis.
 Now we compute the training success rate of TCR and ECR.
-Theorem 3.3.
-Under Assumptions 3.1 and 3.2, the training success rate of TCR in each sample ${\bm{x}}$ is
+Theorem 5.
+Under Assumptions 3 and 4, the training success rate of TCR in each sample $\bm{x}$ is
 […]
-and the training success rate of ECR in each sample ${\bm{x}}$ is
+and the training success rate of ECR is $\forall i\in[n]$,
 […]
-Figure 2: The time consumption during training iterations with different schemes and cluster sizes.
-Remark 3.4.
-In practice, we have constant number of experts [7], i.e., $n=\Theta(1)$, and $C<s$ to save computation cost.
-We explain the benefit of swicthing TCR to ECR during training based on Theorem 3.3 and feature distrution during training.
+Corollary 6.
+In practice, For constant number of experts (Jiang et al.
+2024a), i.e., $n=\Theta(1)$, and $C<s$ to save computation cost.
+We have the following lower bound for capacity $C$ to ensure high training success rate:
+  1. 1.
+    Suppose $q_{i}=\Theta(1)$. Then TCR is much better than ECR, and we only need $C=\Theta(s)$.
+  2. 2.
+    Suppose $\forall i\in[n],sq_{i}\leq C^{*}$ for some $C^{*}>0$. Then ECR is much better than TCR, and we only need $C\geq 2C^{*}$.
+Remark 7.
+We explain the benefit of swicthing TCR to ECR during training based on Theorem 5 and feature distrution during training.
 At the beginning of training, the model seldom learn the task.
-Then the feature of class-irrelevant tokens is nearly isotropy, e.g., uniformly distrbute around the sphere (see Appendix B), leading to $q_{i}=\Theta(1)$.
+Then the feature of class-irrelevant tokens is nearly isotropy, e.g., uniformly distrbute around the sphere (see Appendix), leading to $q_{i}=\Theta(1)$.
 The succed rate of TCR with the form $C/s$ is better than ECR with the form $e^{-s}$.
 Thus we should choose TCR with a large capacity $C=\Theta(s)$ to improve the success rate of training samples.
-After training for some iterations, the experts can roughly distinguish the class-irrelevant and discriminative patterns, leading to $q_{i}\ll 1$ or $sq_{i}\leq C^{*}$ for some $C^{*}>0$ (see Appendix B).
+After training for some iterations, the experts can roughly distinguish the class-irrelevant and discriminative patterns, leading to $q_{i}\ll 1$ or $sq_{i}\leq C^{*}$ for some $C^{*}>0$ (see Appendix).
 Then ECR with success rate nearly $1$ is better than TCR with the form $C/s$ as long as $C\geq 2C^{*}$.
@@
 Indeed, we find that Chowdhury et al.
-[39, the definition of $\ell^{*}$] consider the ECR setting and verify the benefit in sample complexity.
+(2023, the definition of $\ell^{*}$) consider the ECR setting and verify the benefit in sample complexity.
 They assume the maximum number of class-irrelevant patches that are close to class-discriminative patches are bounded, which has similar effect as $C^{*}$ in our scene.
-3.4 Communication Optimization
+Communication Optimization
 The training framework employs the Communication Over Computation (CoC) optimization technique to address performance bottlenecks in LLM training.
@@
 This approach allows for pipeline-style parallel execution and overlapping of computation and communication, significantly enhancing overall efficiency.
-4 Experiments
-4.1 Experimental Setup
-The details of experimental setup including datasets, environment, configuration, and metrics, can be seen in Appendix C.
-4.2 Efficiency Promotion and Memory Footprint Reduction
-Figure 3: (a) The average composition of computation, communication, overlap, and idle with different schemes and cluster sizes. (b) The perplexity during training iterations with different schemes.
-As detailed in Section 3, we consistently use Top-1 routing to ensure the routing implementation aligns with our theoretical framework.
+Experiments
+Experimental Setup
+This study employs the Mixtral 8×7B model, incorporating LocMoE and our proposed approach.
+The Mixtral model, comprising 46.7 billion parameters and utilizing Group Query Attention (GQA), features 32 sparse expert blocks with 8 experts in the MoE Feedforward layer, where each token engages the top 2 experts for processing.
+Given the prevalence of long-text corpora in our application scenarios, we extended the sequence length to 32,768 and implemented tailored parallel strategies for cluster scales of 32N, 64N, and 256N, encompassing tensor, pipeline, data, and expert parallelism, with a consistent global batch size of 128.
+Other details of experimental setup including datasets, environment, configuration, and metrics, can be seen in Appendix.
+Efficiency Promotion and Memory Footprint Reduction
+As detailed in Section Method, we consistently use Top-1 routing to ensure the routing implementation aligns with our theoretical framework.
 The Baseline model utilizes a limited expert capacity mode instead of the groupedGEMM scheme, which avoids token dropping, with the capacity factor set to 1.1.
 LocMoE considers data distribution uniformity and estimates expert capacity using a lower bound formula derived from its theoretical conclusions in the first batch, maintaining it as a constant during subsequent training.
-LocMoE+ fixes the range of score sums, processes hidden states, and calculates current expert capacity.
+Our approach (abbreviate to ”LocMoE+” in figures) fixes the range of score sums, processes hidden states, and calculates current expert capacity.
 The subsequent analysis addresses the training time, convergence, and memory usage efficiency of these schemes on multiple sizes of Ascend clusters.
-Figure 2 illustrates the time consumption of these methods during the first 1000 iterations of training.
+Figure 4: The time consumption during training iterations with different schemes and cluster sizes.
+Figure 3a illustrates the time consumption of these methods during the first 1000 iterations of training.
 Due to initialization and some unstable factors, time consumption is recorded starting from the 5th iteration.
@@
 As iterations increase, LocMoE’s time consumption slightly decreases, particularly in 32N and 64N, consistent with the conclusion that locality loss is effective only when the number of experts is greater than or equal to the number of nodes.
-LocMoE+ incurs slightly higher time consumption than LocMoE due to the computational overhead of token rearrangement.
+Our approach incurs slightly higher time consumption than LocMoE due to the computational overhead of token rearrangement.
 However, as token features converge, the required tokens gradually decrease and stabilize, leading to a decline in time consumption, which remains stable in subsequent training processes.
-Overall, LocMoE+ reduces training time by 2.9% to 13.3% compared to LocMoE, and by 5.4% to 46.6% compared to the Baseline.
+Overall, our approach reduces training time by 2.9% to 13.3% compared to LocMoE, and by 5.4% to 46.6% compared to the Baseline.
 We select 10 iterations at equal intervals from the training iterations to collect data on the time consumption of computation, communication, overlap, and idle periods, as shown in Figure 3(a).
 It is important to note that the data collection operation also introduces some overhead.
-After integrating LocMoE and LocMoE+, the time consumption of each component decreases, with a significantly greater reduction in computation overhead compared to communication overhead.
+After integrating LocMoE and our approach, the time consumption of each component decreases, with a significantly greater reduction in computation overhead compared to communication overhead.
 Additionally, as the cluster size increases, the proportion of computation/communication overlap decreases, and the magnitude of the reduction in computation overhead diminishes.
 Figure 3(b) illustrates perplexity as a measure of convergence.
-The convergence curves of these approaches indicate normal loss convergence, with LocMoE+ not adversely impacting convergence.
+The convergence curves of these approaches indicate normal loss convergence, with our approach not adversely impacting convergence.
+Figure 5: print recorded in one acquisition cycle with different schemes and cluster sizes.
 The proportional time consumption at the operator level is depicted in Figure 6.
 Among the components, AI CORE efficiently executes matrix multiplications and convolutions in AI algorithms; AI VECTOR CORE accelerates vector operations through parallel processing; MIX AIC integrates different types of operators and optimizes for multiple tasks; AI CPU is optimized in hardware and instruction sets to better support AI algorithms.
-LocMoE+ selects fewer tokens, resulting in a 17$\times$ performance improvement in the FFN MatMul operator compared to the Baseline and a 2.6$\times$ improvement compared to LocMoE.
+Our approach selects fewer tokens, resulting in a 17$\times$ performance improvement in the FFN MatMul operator compared to the Baseline and a 2.6$\times$ improvement compared to LocMoE.
 This leads to an overall 2.8$\times$ reduction in the cumulative time consumption of the MatMul operator and a 2.6$\times$ decrease in Cube computing load.
@@
 Figure 6: The distribution of time consumption for operators.
-Figure 5: The footprint recorded in one acquisition cycle with different schemes and cluster sizes.
+Figure 7: The performance on three categories of GDAD.
 We select a single iteration during the stable training period and describe the per-device memory usage (Allocated) using the first 100,000 samples from its memory monitoring, as shown in Figure 5.
-Overall, LocMoE+ achieves memory usage reduction of 4.57% to 16.27% compared to the Baseline and 2.86% to 10.5% compared to LocMoE.
+Overall, our approach achieves memory usage reduction of 4.57% to 16.27% compared to the Baseline and 2.86% to 10.5% compared to LocMoE.
 As cluster size increases, the proportion of computational overhead decreases, and the gap in memory usage narrows.
 Additionally, instantaneous memory peaks gradually disappear, and the fluctuation amplitude of short-term memory also diminishes.
-4.3 The Performance of Downstream Tasks
+The Performance of Downstream Tasks
 To enhance the model’s conversational capabilities and adaptability to downstream task, we fine-tuned the pre-trained models.
-As shown in Figure 6, with sufficient supervised fine-tuning (SFT), LocMoE+ achieves an average improvement of approximately 20.1% in 16 sub-capabilities of Domain Task Capability, which is a portion of General and Domain-specific Assessment Dataset (GDAD), compared to the Baseline, and an increase of about 3.5% compared to LocMoE.
+As shown in Figure 7, with sufficient supervised fine-tuning (SFT), our approach achieves an average improvement of approximately 20.1% in 16 sub-capabilities of Domain Task Capability, which is a portion of General and Domain-specific Assessment Dataset (GDAD), compared to the Baseline, and an increase of about 3.5% compared to LocMoE.
 The Rewriting and Summary capabilities show the highest improvement, with a 28.2% increase compared to the Baseline and a 6.7% increase compared to LocMoE.
-In the 13 tests of Domain Competency Exam, LocMoE+ demonstrates an average improvement of 16% relative to the Baseline and an average increase of approximately 4.8% compared to LocMoE.
+In the 13 tests of Domain Competency Exam, our approach demonstrates an average improvement of 16% relative to the Baseline and an average increase of approximately 4.8% compared to LocMoE.
 The IP Training in the digital communications domain shows the most significant improvement, with a 27.3% increase compared to the Baseline and a 3.0% increase compared to LocMoE.
-Among the 18 sub-capabilities of General Ability, LocMoE+ exhibits an improvement of about 13.9% relative to the Baseline and an average increase of 4.8% compared to LocMoE.
+Among the 18 sub-capabilities of General Ability, our approach exhibits an improvement of about 13.9% relative to the Baseline and an average increase of 4.8% compared to LocMoE.
 The capability of Planning demonstrates the highest improvement, with a 26.8% increase compared to the Baseline and a 2.92% increase compared to LocMoE.
-Figure 6: The performance on three categories of GDAD.
 Table 1 presents the holistic evaluation results for multiple datasets, where GDAD-1 represents Domain Task Capability, and the other metrics follow accordingly.
-Notably, due to the 6:4 ratio of Chinese to English data in our incremental pre-training domain data and the 7:3 ratio in the fine-tuning data, LocMoE+ achieves an improvement of approximately 10.7% compared to the Baseline and 1.2% compared to LocMoE in the C-Eval [40] evaluation, despite the limited data available for training.
+Notably, due to the 6:4 ratio of Chinese to English data in our incremental pre-training domain data and the 7:3 ratio in the fine-tuning data, our approach achieves an improvement of approximately 10.7% compared to the Baseline and 1.2% compared to LocMoE in the C-Eval (Huang et al.
+2024) evaluation, despite the limited data available for training.
 During incremental training and fine-tuning, we incorporated substantial telecommunications domain knowledge, questions, and case studies.
-TeleQnA [41], the first benchmark dataset designed to evaluate the knowledge of LLMs in telecommunications, effectively measures the model’s capabilities in this domain.
-Consequently, LocMoE+ comprehensively surpasses both the Baseline and LocMoE on this specific dataset.
-Table 1: Performance promotion obtained by LocMoE+ on different datasets.
+TeleQnA (Maatouk et al.
+2023), the first benchmark dataset designed to evaluate the knowledge of LLMs in telecommunications, effectively measures the model’s capabilities in this domain.
+Consequently, our approach comprehensively surpasses both the Baseline and LocMoE on this specific dataset.
+Table 1: Performance promotion obtained by our approach on different datasets.
 GDAD
@@
 LocMoE+ 57.4 49.9 74.5 61.5 43.1 68.8
-5 Conclusion
+Conclusion
 In this paper, we propose a novel MoE structure incorporating token feature awareness mechanism to enhance the training efficiency and performance of LLMs.
-Building upon the feature extraction component and locality loss of LocMoE, LocMoE+ introduces a routing strategy that combines TCR and ECR, theoretically demonstrating improved training success rates.
+Building upon the feature extraction component and locality loss of LocMoE, our approach introduces a routing strategy that combines TCR and ECR, theoretically demonstrating improved training success rates.
 By defining the affinity between tokens and experts, expert capacity can be dynamically reduced and further minimized.
 Model training experiments are conducted on Ascend clusters, leveraging the communication optimization characteristics of CoC.
-LocMoE+ achieves performance improvements up to 46.6% (32N) compared to the Baseline and 13.3% (32N) compared to LocMoE, while reducing memory usage by up to 16.27% and 10.5%, respectively.
+Our approach achieves performance improvements up to 46.6% (32N) compared to the Baseline and 13.3% (32N) compared to LocMoE, while reducing memory usage by up to 16.27% and 10.5%, respectively.
 To evaluate model performance, all pre-trained models are fine-tuned using a mixture of domain-specific and general data, and assessed with the open-source datasets C-Eval and TeleQnA, and closed domain benchmark GDAD.
-In downstream tasks, LocMoE+ outperforms the Baseline by 14.1%, 10.7%, and 9.7% on GDAD, C-Eval, and TeleQnA, respectively; it exceeds LocMoE by 4.1%, 1.2%, and 1.7% on these datasets, respectively.
-Both theoretical proofs and experimental results demonstrate the benefits of the LocMoE+ routing strategy for LLM training efficiency.
+In downstream tasks, our approach outperforms the Baseline by 14.1%, 10.7%, and 9.7% on GDAD, C-Eval, and TeleQnA, respectively; it exceeds LocMoE by 4.1%, 1.2%, and 1.7% on these datasets, respectively.
+Both theoretical proofs and experimental results demonstrate the benefits of the proposed routing strategy for LLM training efficiency.
 Future work may explore methods to compress communication data to further reduce communication overhead.
-Appendix A Missing Proofs
-A.1 Auxiurary Results
-Lemma A.1 (Theorem 4 in [42]).
+Appendix A Appendix
+Appendix B Missing Proof
+Auxiurary Results
+Lemma 8 (Theorem 4 in (Chung and Lu 2006)).
 Let $\bm{X}_{1},\dots,\bm{X}_{n}$ be $n$ independent random variables with
-${\mathbb{P}}({\bm{X}}_{i}=1)=p_{i},{\mathbb{P}}({\bm{X}}_{i}=0)=1-p_{i}.$
-We consider the sum ${\bm{X}}=\sum_{i=1}^{n}{\bm{X}}_{i}$, with expectation $\mathbb{E}({\bm{X}})=\sum_{i=1}^{n}p_{i}$. Then we have
-$\mathrm{(Lower\ tail)}\quad{\mathbb{P}}({\bm{X}}\leq\mathbb{E}{\bm{X}}-\lambda)\leq e^{-\frac{\lambda^{2}}{2\mathbb{E}{\bm{X}}}},$ (9)
-$\mathrm{(Upper\ tail)}\quad{\mathbb{P}}({\bm{X}}\geq\mathbb{E}{\bm{X}}+\lambda)\leq e^{-\frac{\lambda^{2}}{2(\mathbb{E}{\bm{X}}+\lambda/3)}}.$
-A.2 Proof of Theorem 3.3
+$\mathcal{P}(\bm{X}_{i}=1)=p_{i},\mathcal{P}(\bm{X}_{i}=0)=1-p_{i}.$
+We consider the sum $\bm{X}=\sum_{i=1}^{n}X_{i}$, with expectation $\mathcal{E}(X)=\sum_{i=1}^{n}p_{i}$. Then we have
+$\mathrm{(Lower\ tail)}\quad\mathcal{P}(\bm{X}\leq\mathcal{E}\bm{X}-\lambda)\leq e^{-\frac{\lambda^{2}}{2\mathcal{E}\bm{X}}},$ (11)
+$\mathrm{(Upper\ tail)}\quad\mathcal{P}(\bm{X}\geq\mathcal{E}\bm{X}+\lambda)\leq e^{-\frac{\lambda^{2}}{2(\mathcal{E}\bm{X}+\lambda/3)}}.$
+Proof of Theorem 5
 Proof.
-1) For the TCR, denote $s_{i}=\left|\left\{t<k:{\bm{x}}_{t}\text{ sent to expert }i,{\bm{x}}_{k}={\bm{o}}_{i}\right\}\right|,\forall i\in[n]$ as the top class-irrelevant token number candidated to the $i$-th expert before the valid token.
-Then by Assumption 3.2, each class-irrelevant token uniformly gives to any expert, leading to $s_{i}|({\bm{x}}_{k}={\bm{o}}_{i})\sim{\mathcal{B}}(k-1,1/n)$ (Binomial distribution), that is,
-${\mathbb{P}}(s_{i}=t|{\bm{x}}_{k}={\bm{o}}_{i})=\begin{pmatrix}k-1\\ t\end{pmatrix}\cdot\left(\frac{1}{n}\right)^{t}\left(1-\frac{1}{n}\right)^{k-1-t},\forall t\in[k-1].$
+1) For the TCR, denote
+$s_{i}=\left|\left\{t<k:\bm{x}_{t}\text{ sent to expert }i,\bm{x}_{k}=\bm{o}_{i}\right\}\right|,\forall i\in[n]$
+as the top class-irrelevant token number candidated to the $i$-th expert before the valid token.
+Then by Assumption 4, each class-irrelevant token uniformly gives to any expert, leading to $s_{i}|(\bm{x}_{k}=\bm{o}_{i})\sim\mathcal{B}(k-1,1/n)$ (Binomial distribution), i.e., $\forall t\in[k-1]$,
+$\mathcal{P}(s_{i}=t|\bm{x}_{k}=\bm{o}_{i})=\begin{pmatrix}k-1\\ t\end{pmatrix}\cdot\left(\frac{1}{n}\right)^{t}\left(1-\frac{1}{n}\right)^{k-1-t}.$
 Then we could derive that
 […]
-$=\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}{\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(C+\sum_{k=C+1}^{s}{\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})\right).$ (10)
-Note that $\mathbb{E}s_{i}=(k-1)/n$. When $k\geq 2nC$, by lower tail bound in Lemma A.1, we get
-${\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})\leq e^{-\frac{(k-1-n(C-1))^{2}}{2(k-1)n}}\leq e^{-\frac{k-1}{8n}}.$
+$=\sum_{i=1}^{n}\mathcal{P}(\bm{o}_{i}\text{ sent to expert }i|\bm{o}_{i}\text{ is in }\bm{x})\cdot\mathcal{P}(\bm{o}_{i}\text{ is in }\bm{x})$
+$=\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$
+$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(C+\sum_{k=C+1}^{s}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\right).$ (12)
+Note that $\mathcal{E}s_{i}=(k-1)/n$. When $k\geq 2nC$, by lower tail bound in Lemma 8, we get
+$\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\leq e^{-\frac{(k-1-n(C-1))^{2}}{2(k-1)n}}\leq e^{-\frac{k-1}{8n}}.$
 Hence, we get the upper bound that
 […]
-$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\sum_{k=2nC+1}^{s}{\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})\right)$
-$\leq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\sum_{k=2nC}^{s-1}e^{-\frac{k}{8n}}\right)\leq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\frac{e^{-\frac{C}{4}}}{1-e^{-\frac{1}{8n}}}\right)$
+(12)${{\leq}}\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$ )
+$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\sum_{k=2nC+1}^{s}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\right)$
+$\leq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\sum_{k=2nC}^{s-1}e^{-\frac{k}{8n}}\right)$
+$\leq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\frac{e^{-\frac{C}{4}}}{1-e^{-\frac{1}{8n}}}\right)$
 $\stackrel{{\scriptstyle(i)}}{{\leq}}\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+(8n+1)e^{-\frac{C}{4}}\right)\leq\frac{10C\sum_{i=1}^{n}p_{i}}{s},$
 where $(i)$ uses the inequality that $e^{-t}\leq 1/(1+t),\forall t\geq 0$.
-Moreover, for $1+\frac{nC}{4}\leq k\leq 1+\frac{nC}{2}$, i.e., $2(k-1)\leq nC\leq 4(k-1)$, by upper tail bound in Lemma A.1, we get
-${\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})=1-{\mathbb{P}}(s_{i}\geq C|{\bm{x}}_{k}={\bm{o}}_{i})\geq 1-e^{-\frac{3(nC-k+1)^{2}}{2n[2(k-1)+nC]}}\geq 1-e^{-\frac{k-1}{4n}}.$
+Moreover, for $1+\frac{nC}{4}\leq k\leq 1+\frac{nC}{2}$, i.e., $2(k-1)\leq nC\leq 4(k-1)$, by upper tail bound in Lemma 8, we get
+$\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$ $=1-\mathcal{P}(s_{i}\geq C|\bm{x}_{k}=\bm{o}_{i})$
+$\geq 1-e^{-\frac{3(nC-k+1)^{2}}{2n[2(k-1)+nC]}}\geq 1-e^{-\frac{k-1}{4n}}.$
 Hence, we get the lower bound that
 […]
-$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\sum_{k=\lceil 1+nC/4\rceil}^{\lfloor 1+nC/2\rfloor}{\mathbb{P}}(s_{i}<C|{\bm{x}}_{k}={\bm{o}}_{i})\right)$
-$\geq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\frac{nC}{4}-1-\sum_{k=\lceil 1+nC/4\rceil}^{\lfloor 1+nC/2\rfloor}e^{-\frac{k-1}{4n}}\right)\geq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\frac{nC}{4}-1-\frac{e^{-\frac{C}{16}}}{1-e^{-\frac{1}{4n}}}\right)$
+(12)${{\geq}}\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$ )
+$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\sum_{k=\lceil 1+nC/4\rceil}^{\lfloor 1+nC/2\rfloor}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\right)$
+$\geq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\frac{nC}{4}-1-\sum_{k=\lceil 1+nC/4\rceil}^{\lfloor 1+nC/2\rfloor}e^{-\frac{k-1}{4n}}\right)$
+$\geq\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\frac{nC}{4}-1-\frac{e^{-\frac{C}{16}}}{1-e^{-\frac{1}{4n}}}\right)$
 $\stackrel{{\scriptstyle(i)}}{{\geq}}\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\frac{nC}{4}-2-(4n+1)e^{-\frac{C}{16}}\right)\geq\frac{C\sum_{i=1}^{n}p_{i}}{5s},$
@@
 2) For the ECR, denote $s_{i}$ as the class-irrelevant token number with the score larger than $\bm{o}_{i}$ for $i$-th expert.
-By Assumption 3.2, we derive that $s_{i}\sim{\mathcal{B}}(s-1,q_{i}),\forall i\in[n]$.
+By Assumption 4, we derive that $s_{i}\sim\mathcal{B}(s-1,q_{i}),\forall i\in[n]$.
 […]
-$=\frac{1}{n}\sum_{i=1}^{n}{\mathbb{P}}(s_{i}\leq C-1,s_{i}\sim{\mathcal{B}}(s-1,q_{i}))$
-If $C-1\leq(s-1)q_{i}/2$, by lower tail bound in Lemma A.1 with $\lambda=(s-1)q_{i}-(C-1)<\mathbb{E}s_{i}$, we obtain that
-${\mathbb{P}}(s_{i}\leq C-1)\leq e^{-\frac{(s-1)q_{i}}{2}\left(1-\frac{C-1}{(s-1)q_{i}}\right)^{2}}\leq e^{-\frac{(s-1)q_{i}}{8}}.$
-If $C\geq 2(s-1)q_{i}$, by upper tail bound in Lemma A.1 with $\lambda=C-(s-1)q_{i}>0$, we obtain that
-${\mathbb{P}}(s_{i}\leq C-1)=1-{\mathbb{P}}(s_{i}\geq C)\geq 1-e^{-\frac{[C-(s-1)q_{i}]^{2}}{2(C+2(s-1)q_{i})/3}}\geq 1-e^{-\frac{3C}{16}}.$
+$=\sum_{i=1}^{n}\mathcal{P}(\text{expert }i\text{ choose }\bm{o}_{i}|\bm{o}_{i}\text{ is in }\bm{x})\mathcal{P}(\bm{o}_{i}\text{ is in }\bm{x})$
+$=\frac{1}{n}\sum_{i=1}^{n}\mathcal{P}(s_{i}\leq C-1,s_{i}\sim\mathcal{B}(s-1,q_{i}))$
+If $C-1\leq(s-1)q_{i}/2$, by lower tail bound in Lemma 8 with $\lambda=(s-1)q_{i}-(C-1)<\mathcal{E}s_{i}$, we obtain that
+$\mathcal{P}(s_{i}\leq C-1)\leq e^{-\frac{(s-1)q_{i}}{2}\left(1-\frac{C-1}{(s-1)q_{i}}\right)^{2}}\leq e^{-\frac{(s-1)q_{i}}{8}}.$
+If $C\geq 2(s-1)q_{i}$, by upper tail bound in Lemma 8 with $\lambda=C-(s-1)q_{i}>0$, we obtain that
+$\mathcal{P}(s_{i}\leq C-1)$ $=1-\mathcal{P}(s_{i}\geq C)$
+$\geq 1-e^{-\frac{[C-(s-1)q_{i}]^{2}}{2(C+2(s-1)q_{i})/3}}\geq 1-e^{-\frac{3C}{16}}.$
 Hence, we conclude Eq. (10). ∎
-Appendix B Token Feature Distribution
+Appendix C Token Feature Distribution
 We also validate the feature distribution before and after MoE training shown in Figure 8.
-We can see before training, all 8192 tokens in one training sample are nearly orthogonal with correlation coefficient near zero, which verifies the isotropy distribution assumption in the first bullet of Remark 3.4.
+We can see before training, all 8192 tokens in one training sample are nearly orthogonal with correlation coefficient near zero, which verifies the isotropy distribution assumption in the first bullet of Remark 7.
 After training, the token features are nearly aligned with correlation coefficien large than $0.8$.
-We can also observe that neighbouring tokens share similar features, and clear block feature behavior, meaning that the token features are relatively separated and the number of tokens in each cluster is bounded, which somehow matches the distribution assumption in the second bullet of Remark 3.4.
+We can also observe that neighbouring tokens share similar features, and clear block feature behavior, meaning that the token features are relatively separated and the number of tokens in each cluster is bounded, which somehow matches the distribution assumption in the second bullet of Remark 7.
 Figure 8: The correlation matrix of one training sample feature before (left) and after (right) training.
-Appendix C Experimental Setup
-C.1 Datasets for Training and Fine-Tuning
+Appendix D Experimental Setup
+Datasets for Training and Fine-Tuning
 The dataset used in this paper is a self-constructed dataset that integrates knowledge from multiple domains, including wireless, data communication, and cloud-core technologies.
@@
 The domain-specific characteristics include domain knowledge understanding, RAG, FunctionCall, information extraction, multi-turn dialogue, reading comprehension, paraphrasing, and intent recognition.
-C.2 Experimental Environment
+Experimental Environment
 The experiments are conducted on a cluster composed of Ascend 910B3 NPUs, divided into three groups: 32 NPUs (hereinafter referred to as 32N, and so on), 64N, and 256N.
@@
 The models in this paper use ModelLink, an LLM training framework based on the Ascend architecture, and run in the torch_npu 5.0.0 environment.
-C.3 Model Configuration
-This paper adopts the Mixtral 8$\times$7B model with an MoE structure, released in December 2023, as the backbone and embeds LocMoE and LocMoE+.
+Model Configuration
+This paper adopts the Mixtral 8$\times$7B model with an MoE structure, released in December 2023, as the backbone and embeds LocMoE and our approach.
 Mixtral has 46.7B parameters and uses the Group Query Attention (GQA) mechanism to compute attention.
@@
 The batch size is set to 128.
-C.4 Evaluation Metrics and Datasets
+Evaluation Metrics and Datasets
 To evaluate model performance, this paper designs a comprehensive metric called the General and Domain-specific Assessment Dataset (GDAD), which consists of three evaluation systems: domain task capability, domain capability certification exam, and general capability.
@@
 For subjective questions, the cosine similarity between the model output and the standard answer is used as the score.
-In addition, this paper also employs C-Eval [40] and TeleQnA [41] to evaluate the model’s Chinese language capability.
+In addition, this paper also employs C-Eval (Huang et al.
+2024) and TeleQnA (Maatouk et al.
+2023) to evaluate the model’s Chinese language capability.

### v3

@@
-Expert-Token Resonance: Redefining MoE Routing through Affinity-Driven Active Selection
+Expert-Token Resonance MoE: Bidirectional Routing with Efficiency Affinity-Driven Active Selection
 Abstract
@@
 However, these architectures grapple with challenges of token distribution imbalance and expert homogenization, impeding optimal semantic generalization.
-We introduce a novel framework that redefines MoE routing through affinity-driven active selection.
-The innovations for the framework encompass: (1) A rigorous formulation of expert-token affinity metrics.
+We propose a novel expert routing framework that incorporates: (1) An efficient routing mechanism with lightweight computation.
 (2) An adaptive bidirectional selection mechanism leveraging resonance between experts and tokens.
-(3) Theoretical derivation and experimental evidence of reduced expert capacity bounds under dynamic token distribution evolution.
+(3) A module that determines the lower bounds of expert capacity based on dynamic token distribution analysis, specifically designed to address drop-and-pad strategies.
 It is also integrated with orthogonal feature extraction module and an optimized loss function for expert localization.
-Our theoretical analysis demonstrates that this approach mitigates expert homogenization while enabling substantial capacity boundary reduction.
-Experimental validation corroborates these findings: it achieves a 40% reduction in token processed by each expert without compromising model convergence or efficacy.
+This framework effectively reduces expert homogeneity while enhancing the performance of the expert selection module.
+Additionally, we introduce a local expert strategy that simultaneously improves load balancing and reduces network communication overhead.
+It achieves a 40% reduction in token processed by each expert without compromising model convergence or efficacy.
 When coupled with communication optimizations, the training efficiency improvements of 5.4% to 46.6% can be observed.
-After supervised fine-tuning, it exhibits performance gains of 9.7% to 14.1% across GDAD, C-Eval, and TeleQnA benchmarks.
+After supervised fine-tuning, it exhibits performance gains of 9.7% to 14.1% across GDAD, GPQA, and TeleQnA benchmarks.
 Introduction
@@
 2022).
-Recently, several LLMs employing MoE structures, such as DeepSeek-MoE (Dai et al.
-2024) and Mixtral (Jiang et al.
+Recently, several LLMs employing MoE structures, such as DeepSeek-V3 (Liu et al.
+2024a) and Mixtral (Jiang et al.
 2024a), have demonstrated outstanding performance on various leaderboards.
@@
 2023).
-DeepSeek-MoE (Dai et al.
-2024) introduces the concept of the vertical-class expert, which is created through fine-grained partitioning, and the shared experts are always activated.
-This approach minimizes the constraints on expert specialization imposed by knowledge hybridity and knowledge redundancy.
 StableMoE (Dai et al.
@@
 This method involves training the routing network independently from the backbone model and utilizing a frozen, distilled routing mechanism to allocate tokens.
+Dynamic-MoE (Huang et al.
+2024a) designs a dynamic routing Mixture-of-Experts (MoE) policy that evaluates the sufficiency of current experts while reducing activated parameters by 90%.
 The characteristics of classical gated routing lead to experts being unable to learn features mastered by other experts.
@@
 This strategy is proven highly effective in achieving expert load balancing and enhancing expert learning outcomes.
+Autonomy-of-Experts models (Lv et al.
+2025) design a novel MoE paradigm in which experts autonomously select themselves to process inputs by aware of its own capacity to effectively process a token.
 The design of routing strategy is crucial to the MoE structure, while not all tokens may be suitable for training (Riquelme et al.
@@
 However, in situations where it is uncertain whether a token contributes to training, there is a risk of discarding class-discriminative samples, potentially compromising the model’s training outcomes.
+DeepSeek-V2 (Liu et al.
+2024a) designs a device-limited routing mechanism to bound MoE-related communication cost.
+DeepSeek-V3 (Liu et al.
+2024a) pioneers an auxiliary-loss-free strategy to minimizes the performance degradation.
+This approach minimizes the constraints on expert specialization imposed by knowledge hybridity and knowledge redundancy.
 XMoE (Yang et al.
@@
   1. 1.
-    A rigorous formulation of expert-token affinity metrics.
-Given that expert selection is related to the angle between tokens and gating weights, the cosine similarity between tokens and gating weights is used to define the affinity score between tokens and experts.
-We provide theoretical analysis of utilizing the affinity score for better guiding expert choice routers to focus on different segmentations of tokens, thus mitigating the homogenization tendency.
-Further, we incorporate affinity scores into the design of our bidirectional selection router.
+    Affinity-based Efficient Expert Routing via GrAP.
+By leveraging cosine similarity between tokens and gating weights to define affinity scores, our router effectively guides experts to focus on distinct token segments, mitigating the expert homogenization problem.
+Meanwhile, the GrAP design reduces computational complexity by a factor of 1/2D to 1/D compared to traditional MLPs (D denotes the dimension of the intermediate hidden layer).
+This integrated approach demonstrates both improved routing effectiveness and substantial computational efficiency.
   2. 2.
@@
 Ultimately, the training efficiency of MoE can be greatly enhanced.
-Expert-token resonance mechanism adopts the state-of-the-art MoE model Mixtral 8$\times$7B as the backbone, and utilizes AscendSpeed and ModelLink libraries for pre-training on Ascend NPU clusters.
+Expert-token resonance mechanism adopts the state-of-the-art MoE model Mixtral 8$\times$7B as the backbone, and utilizes MindSpeed-LLM, MindSpeed, and Megatron-LM (Shoeybi et al.
+2019) libraries for training on Ascend NPU clusters.
+Ascend designs a new computing architecture for LLM training and inference scenarios (Liao et al.
+2021), boasting powerful low-bit computing capabilities.
 Experiments conducted on clusters with 32, 64, and 256 NPUs indicate that our approach improves training efficiency by 5.4% to 46.6% compared to the baseline, and by 2.9% to 13.3% compared to LocMoE.
 Model performance is enhanced by 9.7% to 14.1% compared to the baseline, and by 1.7% to 4.1% compared to LocMoE.
-The rest of this paper is structured as follows: Section Related Work introduces related work on LLMs, MoE, and the Ascend architecture.
-Section Method reviews LocMoE and presents the methods proposed in this paper, along with theoretical evidence.
+The rest of this paper is structured as follows: Section Method presents the methods proposed in this paper, along with theoretical evidence.
 Section Experiments analyzes the experimental results of our approach regarding training efficiency and model performance.
 The final section summarizes the content of this paper and offers an outlook on future improvements.
-Related Work
-LLMs.
-BERT (Kenton and Toutanova 2019) and GPT (Radford et al.
-) in 2018 marked the emergence of the LLM era.
-GPT-3 (Floridi and Chiriatti 2020), with 175 billion parameters, demonstrated impressive capabilities in language understanding and generation in 2020.
-By 2023-2024, LLMs expanded to hundreds of billions to trillions of parameters: GPT-4 (Achiam et al.
-2023) achieved breakthroughs in zero-shot and few-shot learning; PaLM-2 (Anil et al.
-2023) enhanced knowledge acquisition and reasoning; Chinchilla (Hoffmann et al.
-2022) optimized computational resources and model scale; and Claude-v1 (Wei, Haghtalab, and Steinhardt 2024) displayed a safer, human-aligned LLM.
-These LLMs have made significant advancements across domains, paving the way towards artificial general intelligence.
-MoE.
-Despite the outstanding performance of LLMs in language understanding and generation, their high computational cost hinders further model scale expansion.
-To alleviate this issue, researchers have introduced the MoE architecture, which incorporates multiple expert modules and dynamically selects a subset of experts to process inputs, significantly reducing computational overhead while maintaining performance.
-GShard (Lepikhin et al.
-2020) and Switch Transformer (Fedus, Zoph, and Shazeer 2022) employ Top-2 and Top-1 routing strategies, respectively, to mitigate the expert load imbalance problem.
-Hash Layer (Roller et al.
-2021) adopt hashing techniques to simplify the gating mechanism and improve model performance.
-METRO (Bajaj et al.
-2022) introduces a metric learning-based routing mechanism that guides expert selection by learning task-relevant metric functions, enhancing the model’s generalization ability.
-DSelect-k (Hazimeh et al.
-2021) proposes a dynamic selection algorithm based on Markov chains, adaptively choosing the number of experts to maintain performance while reducing training costs.
-Furthermore, DeepSpeed-MoE (Rajbhandari et al.
-2022) optimizes expert parallelism and pipeline parallelism in MoE, further boosting training efficiency.
-These works lay the foundation for the subsequent application of MoE in LLMs.
-Ascend Architecture.
-The AI Core in the Ascend 910B NPU adopts the DaVinci architecture (Tang and Wang 2023), which is responsible for executing computation-intensive operators related to scalars, vectors, and 3D cubes, and is widely used in LLM training.
-Huawei’s proprietary cache coherence protocol interconnect, High Confidence Computing Systems (HCCS) (Xia et al.
-2021), enables high-performance inter-device data communication in multi-card scenarios.
-The Huawei Collective Communication Library (HCCL) (Cao et al.
-2024), designed for the Ascend processor, offers high-performance primitives for single-node multi-card and multi-node multi-card collective communication.
-It implements collective communication functions on PCIe, HCCS, and RoCE high-speed links, facilitating distributed training.
-PyTorch (Paszke et al.
-2019), a widely adopted deep learning framework, is compatible with the Ascend NPU through the Ascend extension for PyTorch (Zhu et al.
-2023).
-AscendSpeed, an acceleration library tailored for Ascend devices, focuses on parallel strategy partitioning for LLMs and aggregates results through collective communication.
-ModelLink, a training framework designed for LLMs, fully considers the characteristics of the Ascend NPU and implements features such as memory optimization, multi-dimensional parallelism, and adaptive data loading, further enhancing training performance.
 Method
-In this section, we present the definition of affinity score with a rigorous formulation of expert-token affinity metrics and the implementation of our affinity-driven active selection routing.
-To reduce the lower bound of expert capacity, we introduce a training strategy for dynamic token distribution evolution under the guiding principle of affinity score.
+In this section, we present the efficient routing mechanism, and our adaptive bidirectional selection mechanism is detailed.
+Then, for traditional drop-and-pad strategies, a dynamic token distribution analysis module that optimizes the lower bounds of expert capacity are displayed.
+Moreover, we also describe the loss for expert load balancing.
 Model Architecture.
@@
 where $n$ is the number of experts, $G(\bm{x}_{t})\colon\mathcal{R}^{d}\to\mathcal{R}^{n}$ is the gating weight vector of experts which maps the tokens of $\bm{x}_{t}$ into the coresponding experts with weights, e.g., $G_{i}(\bm{x})=\text{Softmax}(\bm{W}\bm{x}+\bm{\epsilon})$ where the softmax is applied to each row, and $E_{i}(\bm{x}_{t})\colon\mathcal{R}^{d}\to\mathcal{R}$ is the $i$-th expert network, see (Liu et al.
-2024) for current different router methods.
+2024b) for current different router methods.
 Generally, $n\ll s$, which saves much computation compared to the dense structure.
-Expert-Token Affinity.
-$\bm{W}_{\text{aff}}$ denotes the expert-token affinity matrix. With GrAP as the layer of feature extraction, the formulation of $\bm{W}_{\text{aff}}$ is as followed:
+Cost-Efficient Sparse Expert-Token Affinity.
+$\bm{W}_{\text{aff}}$ denotes the expert-token affinity matrix.
+After processing through the GrAP routing layer, tokens generate a diagonal sparse matrix as shown.
+Compared to the dense matrix produced by traditional routing layers, this reduces the parameter count to 1/D of the original, significantly decreasing the computational overhead of the expert routing layer.
+With GrAP as the layer of feature extraction, the formulation of $\bm{W}_{\text{aff}}$ is as followed:
 $\bm{W}_{\text{aff}}=\begin{pmatrix}\bm{w}_{1}&0&\cdots&0\\ 0&\bm{w}_{2}&\cdots&0\\ \vdots&\vdots&\ddots&\vdots\\ 0&0&\cdots&\bm{w}_{n}\\ \end{pmatrix}$ (2)
@@
 2017) has been proposed.
-Building upon the auxiliary loss, LocMoE introduces a loss bias term based on data locality, represented as $L_{\mathrm{loc}}=\mu\mathrm{KL}(D_{\mathrm{c}}||D_{\mathrm{l}})=-\mu\int{D_{\mathrm{c}}(x)\ln[\frac{D_{\mathrm{l}}(x)}{D_{\mathrm{c}}(x)}]\mathrm{d}x}$, i.e., the Kullback-Leibler (KL) divergence of the current distribution $D_{\mathrm{c}}(x)$ and the fully localized distribution $D_{\mathrm{l}}(x)$.
+Building upon the auxiliary loss, our work introduces a loss bias term based on data locality, represented as $L_{\mathrm{loc}}=\mu\mathrm{KL}(D_{\mathrm{c}}||D_{\mathrm{l}})=-\mu\int{D_{\mathrm{c}}(x)\ln[\frac{D_{\mathrm{l}}(x)}{D_{\mathrm{c}}(x)}]\mathrm{d}x}$, i.e., the Kullback-Leibler (KL) divergence of the current distribution $D_{\mathrm{c}}(x)$ and the fully localized distribution $D_{\mathrm{l}}(x)$.
 This loss term serves as a soft constraint, encouraging tokens to be sent to experts residing on the same node, thereby mitigating the substantial overhead incurred by partial inter-node communication.
 Training Strategy
-The Dynamic Token Distribution Evolution.
-Under the premises of orthogonal gating weights and a data distribution approaching uniformity, LocMoE demonstrates that the expert capacity is closely related to the angle between the gating weights and tokens.
+Token Distribution Dynamics under Expert Routing.
+Under the premises of orthogonal gating weights and a data distribution approaching uniformity, the previous studies demonstrate that the expert capacity is closely related to the angle between the gating weights and tokens.
 For large scale of the activation, the lower bound of expert capacity is proven to exist and is represented as $C_{\mathrm{min}}=\frac{1}{n}\exp\{d\delta_{\mathrm{max}}^{2}/(2-\delta_{\mathrm{max}}^{2})\}$.
@@
 Experimental Setup
-This study employs the Mixtral 8×7B model, incorporating LocMoE and our proposed approach.
+This study employs the Mixtral 8×7B model, incorporating our proposed approach.
 The Mixtral model, comprising 46.7 billion parameters and utilizing Group Query Attention (GQA), features 32 sparse expert blocks with 8 experts in the MoE Feedforward layer, where each token engages the top 2 experts for processing.
 Given the prevalence of long-text corpora in our application scenarios, we extended the sequence length to 32,768 and implemented tailored parallel strategies for cluster scales of 32N, 64N, and 256N, encompassing tensor, pipeline, data, and expert parallelism, with a consistent global batch size of 128.
-Other details of experimental setup including datasets, environment, configuration, and metrics, can be seen in Appendix.
+For the three cluster scales of 32N, 64N, and 256N, the parallel strategies are set as follows: 32N - tensor parallel (TP=4) / pipeline parallel (PP=4) / data parallel (DP=2) / expert parallel (EP=2), 64N - TP=8 / PP=4 / DP=2 / EP=2, and 256N - TP=8 / PP=8 / DP=4 / EP=2.
+Other details of experimental setup including datasets, environment, and metrics, can be seen in Appendix.
 Efficiency Promotion and Memory Footprint Reduction
@@
 Table 1 presents the holistic evaluation results for multiple datasets, where GDAD-1 represents Domain Task Capability, and the other metrics follow accordingly.
-Notably, due to the 6:4 ratio of Chinese to English data in our incremental pre-training domain data and the 7:3 ratio in the fine-tuning data, our approach achieves an improvement of approximately 10.7% compared to the Baseline and 1.2% compared to LocMoE in the C-Eval (Huang et al.
-2024) evaluation, despite the limited data available for training.
+Notably, due to the 6:4 ratio of Chinese to English data in our incremental pre-training domain data and the 7:3 ratio in the fine-tuning data, our approach achieves an improvement of approximately 13.6% compared to the Baseline and 2.8% compared to LocMoE in the GPQA (Rein et al.
+2023) evaluation, despite the limited data available for training.
 During incremental training and fine-tuning, we incorporated substantial telecommunications domain knowledge, questions, and case studies.
@@
 GDAD
-GDAD-1 GDAD-2 GDAD-3 Avg C-Eval TeleQnA
-Baseline 47.8 43.0 65.4 52.8 38.5 62.1
-LocMoE 55.5 47.6 71.1 59.0 42.6 67.6
-LocMoE+ 57.4 49.9 74.5 61.5 43.1 68.8
+GDAD-1 GDAD-2 GDAD-3 Avg GPQA TeleQnA
+Baseline 47.8 43.0 65.4 52.8 29.5 62.1
+LocMoE 55.5 47.6 71.1 59.0 32.6 67.6
+LocMoE+ 57.4 49.9 74.5 61.5 33.5 68.8
 Conclusion
-In this paper, we propose a novel MoE structure incorporating token feature awareness mechanism to enhance the training efficiency and performance of LLMs.
-Building upon the feature extraction component and locality loss of LocMoE, our approach introduces a routing strategy that combines TCR and ECR, theoretically demonstrating improved training success rates.
-By defining the affinity between tokens and experts, expert capacity can be dynamically reduced and further minimized.
-Model training experiments are conducted on Ascend clusters, leveraging the communication optimization characteristics of CoC.
+In this paper, we propose a novel expert routing framework that enhances MoE efficiency through three key innovations: an efficient routing mechanism with lightweight computation, a bidirectional expert-token resonance selection mechanism, which combined ECR and TCR, and a dynamic capacity bounds module.
+The framework integrates orthogonal feature extraction and optimized expert localization loss, effectively addressing expert homogeneity while improving routing performance.
+Our local expert strategy demonstrates advantages in both load balancing and communication efficiency.
+Experimental results validate the effectiveness of the proposed framework across multiple benchmarks.
 Our approach achieves performance improvements up to 46.6% (32N) compared to the Baseline and 13.3% (32N) compared to LocMoE, while reducing memory usage by up to 16.27% and 10.5%, respectively.
-To evaluate model performance, all pre-trained models are fine-tuned using a mixture of domain-specific and general data, and assessed with the open-source datasets C-Eval and TeleQnA, and closed domain benchmark GDAD.
-In downstream tasks, our approach outperforms the Baseline by 14.1%, 10.7%, and 9.7% on GDAD, C-Eval, and TeleQnA, respectively; it exceeds LocMoE by 4.1%, 1.2%, and 1.7% on these datasets, respectively.
-Both theoretical proofs and experimental results demonstrate the benefits of the proposed routing strategy for LLM training efficiency.
+To evaluate model performance, all models are evaluated with the open-source datasets GPQA and TeleQnA, and closed domain benchmark GDAD.
+In downstream tasks, our approach outperforms the Baseline by 14.1%, 13.6%, and 9.7% on GDAD, GPQA, and TeleQnA, respectively.
 Future work may explore methods to compress communication data to further reduce communication overhead.
-Appendix A Appendix
-Appendix B Missing Proof
+Appendix
+Missing Proof
 Auxiurary Results
@@
 Let $\bm{X}_{1},\dots,\bm{X}_{n}$ be $n$ independent random variables with
-$\mathcal{P}(\bm{X}_{i}=1)=p_{i},\mathcal{P}(\bm{X}_{i}=0)=1-p_{i}.$
+$\mathcal{P}(\bm{X}_{i}=1)=p_{i},\mathcal{P}(\bm{X}_{i}=0)=1-p_{i}.$ (11)
 We consider the sum $\bm{X}=\sum_{i=1}^{n}X_{i}$, with expectation $\mathcal{E}(X)=\sum_{i=1}^{n}p_{i}$. Then we have
@@
 1) For the TCR, denote
-$s_{i}=\left|\left\{t<k:\bm{x}_{t}\text{ sent to expert }i,\bm{x}_{k}=\bm{o}_{i}\right\}\right|,\forall i\in[n]$
+$s_{i}=\left|\left\{t<k:\bm{x}_{t}\text{ sent to expert }i,\bm{x}_{k}=\bm{o}_{i}\right\}\right|,\forall i\in[n]$ (13)
 as the top class-irrelevant token number candidated to the $i$-th expert before the valid token.
 Then by Assumption 4, each class-irrelevant token uniformly gives to any expert, leading to $s_{i}|(\bm{x}_{k}=\bm{o}_{i})\sim\mathcal{B}(k-1,1/n)$ (Binomial distribution), i.e., $\forall t\in[k-1]$,
-$\mathcal{P}(s_{i}=t|\bm{x}_{k}=\bm{o}_{i})=\begin{pmatrix}k-1\\ t\end{pmatrix}\cdot\left(\frac{1}{n}\right)^{t}\left(1-\frac{1}{n}\right)^{k-1-t}.$
+$\mathcal{P}(s_{i}=t|\bm{x}_{k}=\bm{o}_{i})=\begin{pmatrix}k-1\\ t\end{pmatrix}\cdot\left(\frac{1}{n}\right)^{t}\left(1-\frac{1}{n}\right)^{k-1-t}.$ (14)
 Then we could derive that
@@
 $=\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$
-$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(C+\sum_{k=C+1}^{s}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\right).$ (12)
+$=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(C+\sum_{k=C+1}^{s}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\right).$
 Note that $\mathcal{E}s_{i}=(k-1)/n$. When $k\geq 2nC$, by lower tail bound in Lemma 8, we get
-$\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\leq e^{-\frac{(k-1-n(C-1))^{2}}{2(k-1)n}}\leq e^{-\frac{k-1}{8n}}.$
+$\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\leq e^{-\frac{(k-1-n(C-1))^{2}}{2(k-1)n}}\leq e^{-\frac{k-1}{8n}}.$ (15)
 Hence, we get the upper bound that
 […]
-(12)${{\leq}}\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$ )
+(Proof.)${{\leq}}\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$ )
 $=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\sum_{k=2nC+1}^{s}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\right)$
@@
 […]
-(12)${{\geq}}\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$ )
+(Proof.)${{\geq}}\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$ )
 $=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\sum_{k=\lceil 1+nC/4\rceil}^{\lfloor 1+nC/2\rfloor}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\right)$
@@
 If $C-1\leq(s-1)q_{i}/2$, by lower tail bound in Lemma 8 with $\lambda=(s-1)q_{i}-(C-1)<\mathcal{E}s_{i}$, we obtain that
-$\mathcal{P}(s_{i}\leq C-1)\leq e^{-\frac{(s-1)q_{i}}{2}\left(1-\frac{C-1}{(s-1)q_{i}}\right)^{2}}\leq e^{-\frac{(s-1)q_{i}}{8}}.$
+$\mathcal{P}(s_{i}\leq C-1)\leq e^{-\frac{(s-1)q_{i}}{2}\left(1-\frac{C-1}{(s-1)q_{i}}\right)^{2}}\leq e^{-\frac{(s-1)q_{i}}{8}}.$ (16)
 If $C\geq 2(s-1)q_{i}$, by upper tail bound in Lemma 8 with $\lambda=C-(s-1)q_{i}>0$, we obtain that
@@
 Hence, we conclude Eq. (10). ∎
-Appendix C Token Feature Distribution
+Token Feature Distribution
 We also validate the feature distribution before and after MoE training shown in Figure 8.
@@
 Figure 8: The correlation matrix of one training sample feature before (left) and after (right) training.
-Appendix D Experimental Setup
+Experimental Setup
 Datasets for Training and Fine-Tuning
@@
 In addition, this paper also employs C-Eval (Huang et al.
-2024) and TeleQnA (Maatouk et al.
+2024b) and TeleQnA (Maatouk et al.
 2023) to evaluate the model’s Chinese language capability.

### v4

@@
 Abstract
-Mixture-of-Experts (MoE) architectures have emerged as a paradigm-shifting approach for large language models (LLMs), offering unprecedented computational efficiency.
-However, these architectures grapple with challenges of token distribution imbalance and expert homogenization, impeding optimal semantic generalization.
-We propose a novel expert routing framework that incorporates: (1) An efficient routing mechanism with lightweight computation.
-(2) An adaptive bidirectional selection mechanism leveraging resonance between experts and tokens.
-(3) A module that determines the lower bounds of expert capacity based on dynamic token distribution analysis, specifically designed to address drop-and-pad strategies.
-It is also integrated with orthogonal feature extraction module and an optimized loss function for expert localization.
-This framework effectively reduces expert homogeneity while enhancing the performance of the expert selection module.
-Additionally, we introduce a local expert strategy that simultaneously improves load balancing and reduces network communication overhead.
-It achieves a 40% reduction in token processed by each expert without compromising model convergence or efficacy.
-When coupled with communication optimizations, the training efficiency improvements of 5.4% to 46.6% can be observed.
-After supervised fine-tuning, it exhibits performance gains of 9.7% to 14.1% across GDAD, GPQA, and TeleQnA benchmarks.
+Mixture-of-Experts (MoE) architectures enable efficient scaling of large language models by activating only a subset of parameters per input.
+However, existing MoE models suffer from two critical limitations: (1) inefficient token-to-expert routing that causes excessive communication overhead, and (2) expert homogenization that leads to redundant computations.
+Current approaches address these challenges separately, failing to achieve simultaneous improvements in both training efficiency and model performance.
+We present Expert-Token Resonance (ETR), a theoretically-grounded bidirectional routing mechanism that fundamentally reimagines expert-token interactions in MoE architectures.
+Our key insight is that optimal routing requires adaptive coordination between token-choice routing (TCR) during early training phases and expert-choice routing (ECR) in later stages.
+We prove that this dynamic approach maximizes training success rate—the probability of correct token-expert assignments—while reducing the expert capacity lower bound by up to 40%.
+ETR incorporates three technical innovations: (1) an affinity-based routing architecture using Grouped Average Pooling (GrAP) that reduces computational complexity from O(d²) to O(d²/D) while maintaining orthogonality to prevent expert homogenization; (2) a bidirectional selection mechanism that enables both tokens and experts to actively participate in the routing process based on cosine similarity scores; and (3) an adaptive capacity strategy that dynamically adjusts expert bounds based on training progress, eliminating communication bubbles in All-to-All operations.
+Extensive experiments on Ascend NPU clusters demonstrate that ETR achieves 5.4%-46.6% improvements in end-to-end training efficiency compared to baseline MoE implementations, with 9.7%-14.5% performance gains across GDAD, GPQA, HumanEval, and TeleQnA benchmarks.
+These results establish ETR as the efficient MoE routing approach to deliver substantial improvements in both computational efficiency and model quality, enabling practical deployment of larger sparse models previously constrained by communication bottlenecks.
 Introduction
-Large language models (LLMs) have shown exceptional proficiency in understanding deep structures and complex semantic relationships within language (Zhao et al.
+Large language models (LLMs) have demonstrated remarkable capabilities in understanding complex semantic relationships and generating coherent text (Zhao et al.
 2023).
-As these models scale up, their capabilities in language generation and logical comprehension are enhanced, but this comes at the cost of significant computational, communication, and storage demands (Jiang et al.
+However, as model parameters scale to billions or even trillions, the computational and communication costs grow prohibitively (Jiang et al.
 2024b).
-To scale models efficiently without disproportionately increasing computational costs, researchers have incorporated the Mixture-of-Experts (MoE) architecture into LLMs (Lepikhin et al.
-2020).
-The MoE framework integrates multiple experts within the model, each tasked with processing specific types of inputs (Fedus, Zoph, and Shazeer 2022).
-For a given input, only a subset of experts is activated, allowing for more efficient use of computational resources (Du et al.
+The Mixture-of-Experts (MoE) architecture offers a promising solution by activating only a subset of model parameters for each input, enabling efficient model scaling without proportional increases in computational demands (Lepikhin et al.
+2020; Fedus, Zoph, and Shazeer 2022).
+Recent MoE-based LLMs, such as DeepSeek-V3 (Liu et al.
+2024a) and Mixtral (Jiang et al.
+2024a), have achieved state-of-the-art performance across various benchmarks.
+Despite these successes, MoE models face two fundamental challenges that limit their practical deployment and effectiveness.
+First, the routing efficiency problem: current routing mechanisms suffer from suboptimal token-to-expert assignments, leading to significant computational waste through excessive padding in All-to-All communications and underutilized expert capacity.
+Second, the expert homogenization problem: existing routing strategies often fail to maintain expert specialization, resulting in redundant computations where multiple experts learn similar representations.
+These challenges are particularly pronounced during early training phases when routing networks lack sophisticated assignment capabilities.
+Previous approaches have attempted to address these issues separately.
+Load balancing techniques (Zhou et al.
+2022b; Dai et al.
+2022) focus on distributing tokens evenly across experts but often sacrifice routing quality.
+Expert specialization methods (Xie et al.
+2024; Li et al.
+2024) aim to maintain diversity but typically incur additional computational overhead.
+Critically, no existing work has demonstrated how to simultaneously improve both training efficiency and model performance through a unified routing mechanism.
+This paper introduces Expert-Token Resonance (ETR), a novel bidirectional routing strategy that fundamentally rethinks how tokens and experts interact in MoE architectures.
+Our key insight is that optimal routing requires not only tokens selecting appropriate experts (token-choice routing, TCR) but also experts actively selecting their most relevant tokens (expert-choice routing, ECR).
+We theoretically prove that TCR achieves higher training success rates during early training when routing networks need refinement, while ECR excels in later stages when expert specialization becomes critical.
+By adaptively combining both paradigms, ETR achieves what previous methods could not: simultaneous improvements in training efficiency and model performance.
+Specifically, our contributions are threefold:
+  1. 1.
+    A theoretically-grounded bidirectional routing mechanism that dynamically balances token-choice and expert-choice routing based on training progress.
+We prove that this approach maximizes the training success rate—the probability of correct token-expert assignments—throughout the entire training process while reducing the required expert capacity lower bound by up to 40% compared to conventional methods.
+  2. 2.
+    An efficient affinity-based routing architecture leveraging Grouped Average Pooling (GrAP) layers that reduces computational complexity from $O(d^{2})$ to $O(d^{2}/D)$ compared to traditional MLP routers, where $d$ is the hidden dimension and $D$ is the grouping factor.
+The orthogonality properties of GrAP naturally prevent expert homogenization while enabling precise affinity computations through cosine similarity.
+  3. 3.
+    Comprehensive empirical validation on Ascend NPU clusters demonstrating that ETR improves end-to-end training efficiency by 5.4%-46.6% compared to baseline MoE implementations and 2.9%-13.3% compared to state-of-the-art LocMoE, while simultaneously enhancing downstream task performance by 9.7%-14.5% and 1.7%-4.1%, respectively.
+The significance of this work extends beyond incremental improvements.
+By addressing the fundamental trade-off between routing quality and computational efficiency, ETR enables practical deployment of larger MoE models that were previously constrained by communication bottlenecks.
+Our theoretical analysis provides new insights into the dynamics of expert-token interactions, opening avenues for future research in sparse model architectures.
+Related Work
+Routing Mechanisms
+Routing design fundamentally determines MoE performance.
+Token-choice routing (TCR) (Shazeer et al.
+2017a) allows tokens to select experts but suffers from load imbalance.
+Expert-choice routing (ECR) (Zhou et al.
+2022a) ensures balanced loads by having experts select tokens, risking the loss of critical tokens.
+Recent dynamic routing strategies (Huang et al.
+2024) adapt expert allocation to input complexity but remain unidirectional—either tokens choose experts or vice versa.
+This fundamental limitation prevents simultaneous optimization of routing quality and load balance.
+Expert Specialization
+Expert homogenization critically limits MoE performance.
+Prior work addressed this through fine-grained expert segmentation (Dai et al.
+2024), competitive mechanisms (Pham et al.
+2024), and theoretical analyses linking data clustering to specialization (Chen et al.
 2022).
-Recently, several LLMs employing MoE structures, such as DeepSeek-V3 (Liu et al.
-2024a) and Mixtral (Jiang et al.
-2024a), have demonstrated outstanding performance on various leaderboards.
-Despite the efficiency benefits of MoE in scaling model sizes, it introduces several new challenges and drawbacks (Shazeer et al.
-2017).
-The conventional MoE model’s convergence and the experts’ generalization capabilities are heavily dependent on the design of the routing strategy, which easily leads to an imbalanced ”winner-takes-all” phenomenon among experts.
-The imbalance between excessively ”developed” experts and those lacking adequate training may compromise or even nullify the intended functionality of routing strategies.
-Recent studies address these challenges from multiple perspectives (Li et al.
-2023).
-StableMoE (Dai et al.
-2022) proposes a two-stage training approach to address the issue of routing fluctuation.
-This method involves training the routing network independently from the backbone model and utilizing a frozen, distilled routing mechanism to allocate tokens.
-Dynamic-MoE (Huang et al.
-2024a) designs a dynamic routing Mixture-of-Experts (MoE) policy that evaluates the sufficiency of current experts while reducing activated parameters by 90%.
-The characteristics of classical gated routing lead to experts being unable to learn features mastered by other experts.
-To address this, MoDE (Xie et al.
-2024) proposes moderate distillation between experts to mitigate the generalization problems caused by narrow learning paths.
-DYNMoE (Guo et al.
-2024) introduces a unique gated routing mechanism capable of adaptively determining the number of activated experts through trainable expert thresholds, even allowing for the addition or removal of experts.
-In addition to the classical token choice scenario, previous researches also propose work utilizing expert choice (EC).
-Google Brain introduces the EC routing algorithm (Zhou et al.
-2022), which assigns experts with predetermined buffer capacities to the Top-k tokens to ensure load balance.
-The Brainformer (Zhou et al.
-2023) also adopts this routing strategy, constructing a trainable gating matrix to project the input feature space onto scores corresponding to each expert.
-Then, each token is routed to the Top-k experts.
-This strategy is proven highly effective in achieving expert load balancing and enhancing expert learning outcomes.
-Autonomy-of-Experts models (Lv et al.
-2025) design a novel MoE paradigm in which experts autonomously select themselves to process inputs by aware of its own capacity to effectively process a token.
-The design of routing strategy is crucial to the MoE structure, while not all tokens may be suitable for training (Riquelme et al.
-2021).
-In addition to data preprocessing techniques such as dataset cleaning and deduplication, previous studies have also considered how to discard certain tokens within the model.
-Early work introduced the concept of expert capacity (Lepikhin et al.
-), which refers to the maximum number of tokens each expert can process at once.
-Tokens exceeding this capacity are discarded.
-Expert capacity helps to ensure load balance among experts while facilitating All-to-All communication implementation.
-However, in situations where it is uncertain whether a token contributes to training, there is a risk of discarding class-discriminative samples, potentially compromising the model’s training outcomes.
-DeepSeek-V2 (Liu et al.
-2024a) designs a device-limited routing mechanism to bound MoE-related communication cost.
-DeepSeek-V3 (Liu et al.
-2024a) pioneers an auxiliary-loss-free strategy to minimizes the performance degradation.
-This approach minimizes the constraints on expert specialization imposed by knowledge hybridity and knowledge redundancy.
-XMoE (Yang et al.
-2024) achieves more precise router by implementing a threshold-based approach.
-If a token reaches the specified threshold, it is processed exclusively by a single expert while being discarded by other experts within the Top-k selection.
-This method allows for more nuanced token selection and processing.
-LocMoE (Li et al.
-2024) leverages orthogonal routing weights to prevent token homogenization across different expert networks and introduces the Grouped Average Pooling (GrAP) layer (Wang, Zhang, and Du 2023) for token feature extraction.
-Under these conditions, LocMoE also provides the theoretical proof for the lower bound of expert capacity.
-In this paper, we propose expert-token resonance, a mechanism consisting of an expert-token bidirectional selection router and the adaptive expert capacity strategy.
-The primary contributions of this paper are as follows:
-  1. 1.
-    Affinity-based Efficient Expert Routing via GrAP.
-By leveraging cosine similarity between tokens and gating weights to define affinity scores, our router effectively guides experts to focus on distinct token segments, mitigating the expert homogenization problem.
-Meanwhile, the GrAP design reduces computational complexity by a factor of 1/2D to 1/D compared to traditional MLPs (D denotes the dimension of the intermediate hidden layer).
-This integrated approach demonstrates both improved routing effectiveness and substantial computational efficiency.
-  2. 2.
-    Expert-token Bidirectional Selection.
-By integrating the concepts of expert choice router (ECR) and token choice router (TCR), we propose the adaptive bidirectional selection mechanism.
-Contrast to conventional router, the bidirectional selection router allows MoE to enhance the training success rate while considering expert capacity constraints.
-Its effectiveness has been theoretically validated.
-  3. 3.
-    The Adaptive Expert Capacity Bound.
-Setting an adaptive affinity threshold allows the lower bound of expert capacity to be significantly reduced.
-As training iterations increase, the information density of token features grows, causing the expert capacity to initially decrease and then stabilize.
-Ultimately, the training efficiency of MoE can be greatly enhanced.
-Expert-token resonance mechanism adopts the state-of-the-art MoE model Mixtral 8$\times$7B as the backbone, and utilizes MindSpeed-LLM, MindSpeed, and Megatron-LM (Shoeybi et al.
-2019) libraries for training on Ascend NPU clusters.
-Ascend designs a new computing architecture for LLM training and inference scenarios (Liao et al.
-2021), boasting powerful low-bit computing capabilities.
-Experiments conducted on clusters with 32, 64, and 256 NPUs indicate that our approach improves training efficiency by 5.4% to 46.6% compared to the baseline, and by 2.9% to 13.3% compared to LocMoE.
-Model performance is enhanced by 9.7% to 14.1% compared to the baseline, and by 1.7% to 4.1% compared to LocMoE.
-The rest of this paper is structured as follows: Section Method presents the methods proposed in this paper, along with theoretical evidence.
-Section Experiments analyzes the experimental results of our approach regarding training efficiency and model performance.
-The final section summarizes the content of this paper and offers an outlook on future improvements.
+However, these approaches require auxiliary losses or complex training procedures.
+System Optimization
+System-level MoE optimizations focus on hardware utilization through block-sparse operations (Gale et al.
+2022) and communication reduction via expert sharding (Balmau et al.
+2025) or sequence migration (Chen et al.
+2024).
+While these approaches improve specific bottlenecks, they overlook communication bubbles in All-to-All operations—a critical inefficiency.
+Positioning of This Work
+Recent surveys (Cai et al.
+2024) reveal that existing MoE methods typically prioritize either routing quality or computational efficiency.
+Our ETR framework addresses previously overlooked communication inefficiencies and training bottleneck, suggesting potential for more efficient MoE deployment at scale.
 Method
@@
 Moreover, we also describe the loss for expert load balancing.
+Figure 1: The illustrative diagram of GrAP.
 Model Architecture.
@@
 Compared to the dense matrix produced by traditional routing layers, this reduces the parameter count to 1/D of the original, significantly decreasing the computational overhead of the expert routing layer.
+As shown in Figure 1, GrAP performs average pooling with group segmented by the number of experts.
 With GrAP as the layer of feature extraction, the formulation of $\bm{W}_{\text{aff}}$ is as followed:
@@
 2022; Zhou et al.
-2022) and ECR.
+2022b) and ECR.
 As the name suggested, TCR lets each token choose its top-scored experts, and ECR lets each expert choose its top-scored tokens.
@@
 $~~\tilde{I}_{tk}\in[n],\forall t\in[s],k\in[\ell].$
-and then the expert to choose its $-\ell$ tokens where $\ell$ is determined by a threshold of the sum of affinity scores:
+and then the expert to choose its Top-$\ell$ tokens where $\ell$ is determined by a threshold of the sum of affinity scores:
 $\left(I_{1i},\dots,I_{Ci}\right)=\text{Bottom-}C\left(\left\{t\in[s]:\exists j\in[\ell],\tilde{I}_{tj}=i\right\}\right),$ (6)
@@
 To mitigate this skewness in token allocation, the auxiliary loss (Shazeer et al.
-2017) has been proposed.
+2017b) has been proposed.
 Building upon the auxiliary loss, our work introduces a loss bias term based on data locality, represented as $L_{\mathrm{loc}}=\mu\mathrm{KL}(D_{\mathrm{c}}||D_{\mathrm{l}})=-\mu\int{D_{\mathrm{c}}(x)\ln[\frac{D_{\mathrm{l}}(x)}{D_{\mathrm{c}}(x)}]\mathrm{d}x}$, i.e., the Kullback-Leibler (KL) divergence of the current distribution $D_{\mathrm{c}}(x)$ and the fully localized distribution $D_{\mathrm{l}}(x)$.
@@
 Token Distribution Dynamics under Expert Routing.
-Under the premises of orthogonal gating weights and a data distribution approaching uniformity, the previous studies demonstrate that the expert capacity is closely related to the angle between the gating weights and tokens.
-For large scale of the activation, the lower bound of expert capacity is proven to exist and is represented as $C_{\mathrm{min}}=\frac{1}{n}\exp\{d\delta_{\mathrm{max}}^{2}/(2-\delta_{\mathrm{max}}^{2})\}$.
-The hybrid TCR+ECR bidirectional selection routing, introduced in the model structure, is exemplified in the figure.
-If the feature fragment corresponding to the $k$-th dimension of the gating weight for a particular token is more prominent, then that token will be routed to the $k$-th expert.
-If among all tokens routed to the $k$-th expert, there is a certain probability of the presence of class-discriminative tokens, then the capacity $C$ must be set to a larger value to ensure the inclusion of sufficient class-discriminative tokens.
-The router proposed in this paper is a hybrid of TCR and ECR modes.
-After determining the expert to which a token will be routed, scores are calculated for the tokens assigned to each expert, and a $-\ell$ selection is performed, where $\ell^{*}$ is determined by a threshold of the sum of scores.
-Subsequent theoretical analysis will demonstrate the effectiveness of this hybrid routing scheme.
-Figure 3: (a) The average composition of computation, communication, overlap, and idle with different schemes and cluster sizes. (b) The perplexity during training iterations with different schemes.
-Theoretical Explanation
-To explain the motivation of our method, we show some theoretical insights in this section. Our theoretical analyis is bulit on Chowdhury et al. (2023), where they make the following data assumption:
+The proposed hybrid TCR+ECR bidirectional routing mechanism operates through a two-stage process: tokens are initially assigned to experts based on the similarity between their feature fragments and the corresponding gating weights, followed by each expert selecting the Top-$\ell$ most relevant tokens through a scoring mechanism.
+This hybrid approach dynamically determines the number of tokens processed by each expert via an adaptive threshold, thereby ensuring the retention of class-discriminative tokens while optimizing computational efficiency.
+Dynamic Lower Bound Module for Expert Capacity in ETR
+To explain the motivation of our method, we show some theoretical insights in this section.
 Assumption 1 (data assumption).
-Each input $\bm{x}\in\mathcal{R}^{sd}$ with $s$ tokens is comprised of one class-discriminative pattern $\bm{o}_{1},\dots,\bm{o}_{n}\in\mathcal{R}^{d}$, with each decides the label in $[n]$, and $s-1$ class-irrelevant patterns $\bm{r}\sim\mathcal{N}$ for certain distribution $\mathcal{N}$.
+Each input $\bm{x}\in\mathcal{R}^{sd}$ with $s$ tokens is comprised of one class-discriminative pattern $\bm{o}_{1},\dots,\bm{o}_{n}\in\mathcal{R}^{d}$, with each decides the label in $[n]$, and s1s-1 class-irrelevant patterns $\bm{r}\sim\mathcal{N}$ for certain distribution $\mathcal{N}$.
 For example, $\bm{x}=(\bm{r}_{1},\bm{r}_{2},\bm{o}_{1},\bm{r}_{3},\dots,\bm{r}_{s-1})$ has label $1$, where $\bm{r}_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N},\forall i\in[s-1]$.
 Based on Assumption 1, Chowdhury et al. (2023) demonstrated that the training of MoE go through two phases:
-Phase 1: Router training (Chowdhury et al.
-2023, Lemma 4.1 and Assumption 4.4), which makes class-discriminative patterns all to the corresponding expert.
-This process ensures that each expert only receives the class-discriminative tokens related to the specific class.
-Phase 2: Expert training (Chowdhury et al.
-2023, Theorem 4.2 and Theorme 4.5), which makes each expert learn to predict the label based on its class-discriminative inputs from Phase 1.
-This process is designed to establish each expert’s ability to handle and solve problems.
-Hence, the traning of an input in the current step is valid if the class-discriminative patterns is correctly dispatched.
+Phase 1: Router training This process ensures that each expert only receives the class-discriminative tokens related to the specific class.
+Phase 2: Expert training This process is designed to establish each expert’s ability to handle and solve problems.
 To quantitatively measure the difference between TCR and ECR, we define training success rate of input motivated by the training process of MoE.
@@
 We assume the location and feature of class-discriminative pattern is uniformly distribute in $[s]$ and $[n]$, i.e.,
-$i\sim\mathrm{Unif}([s]),\bm{x}_{i}\sim\mathrm{Unif}\left(\{\bm{o}_{1},\dots,\bm{o}_{n}\}\right).$
+$i\sim\mathrm{Unif}([s]),\bm{x}_{i}\sim\mathrm{Unif}\left(\{\bm{o}_{1},\dots,\bm{o}_{n}\}\right).$ (7)
 We also assume that $\forall i\in[n],\bm{o}_{i}$ should be sent to the $i$-th expert, and define the true positive probability in token choice setting is no worse than the uniform dispatch as below
-$\mathcal{P}(\delta_{\bm{o}_{i},i}\geq\delta_{\bm{x}_{j},i},\forall j\in[s])=p_{i}\geq 1/n,\forall i\in[n].$
+$\mathcal{P}(\delta_{\bm{o}_{i},i}\geq\delta_{\bm{x}_{j},i},\forall j\in[s])=p_{i}\geq 1/n,\forall i\in[n].$ (8)
 Assumption 4 (class-irrelevant).
@@
 which measures the possibility that expert $i$ chooses the wrong token $\bm{r}$ instead of the correct token $\bm{o}_{i}$.
-Assumption 3 assumes the valid token is uniformly distributed in training samples due to the massive amounts of data nowadays.
-Assumption 4 assumes the invalid tokens can be uniformly dispatched to experts since the invalid tokens do not provide supervised signal to router and experts in training.
-We consider such uniform settings are common assumptions in theoretical analyis.
-Now we compute the training success rate of TCR and ECR.
 Theorem 5.
 Under Assumptions 3 and 4, the training success rate of TCR in each sample $\bm{x}$ is
-[…]
+$\mathcal{P}(\text{TCR succeed})=\Theta\big(C\sum_{i=1}^{n}p_{i}/s\big),$ (11)
 and the training success rate of ECR is $\forall i\in[n]$,
-[…]
+$\mathcal{P}(\text{ECR succeed})\begin{cases}\leq\frac{1}{n}\sum_{i=1}^{n}e^{-\frac{(s-1)q_{i}}{8}},&C\leq(s-1)q_{i}/2,\\ \geq 1-e^{-3C/16},&C\geq 2sq_{i}.\end{cases}$ (12)
 Corollary 6.
@@
 Remark 7.
-We explain the benefit of swicthing TCR to ECR during training based on Theorem 5 and feature distrution during training.
-At the beginning of training, the model seldom learn the task.
-Then the feature of class-irrelevant tokens is nearly isotropy, e.g., uniformly distrbute around the sphere (see Appendix), leading to $q_{i}=\Theta(1)$.
-The succed rate of TCR with the form $C/s$ is better than ECR with the form $e^{-s}$.
-Thus we should choose TCR with a large capacity $C=\Theta(s)$ to improve the success rate of training samples.
-After training for some iterations, the experts can roughly distinguish the class-irrelevant and discriminative patterns, leading to $q_{i}\ll 1$ or $sq_{i}\leq C^{*}$ for some $C^{*}>0$ (see Appendix).
-Then ECR with success rate nearly $1$ is better than TCR with the form $C/s$ as long as $C\geq 2C^{*}$.
-Thus we should choose ECR with a small capacity $C=\Theta(1)$ to improve the success rate of training samples.
-Indeed, we find that Chowdhury et al.
-(2023, the definition of $\ell^{*}$) consider the ECR setting and verify the benefit in sample complexity.
-They assume the maximum number of class-irrelevant patches that are close to class-discriminative patches are bounded, which has similar effect as $C^{*}$ in our scene.
+Under the assumptions of orthogonal gating weights and uniform data distribution, we establish that the lower bound of expert capacity is given by $C_{\mathrm{min}}=\frac{1}{n}\exp\{d\delta_{\mathrm{max}}^{2}/(2-\delta_{\mathrm{max}}^{2})\}$, where the capacity is intrinsically linked to the angular relationship between gating weights and tokens.
+Building on Theorem 5 and the evolution of feature distributions during training, we demonstrate the optimality of transitioning from TCR to ECR.
+In early training stages, when class-irrelevant tokens exhibit near-isotropic distribution with $q_{i}=\Theta(1)$, TCR achieves superior training success rates of $C/s$ compared to ECR’s exponentially decaying rate of $e^{-s}$, necessitating a larger expert capacity of $C=\Theta(s)$.
+As training progresses and experts develop discriminative capabilities, the distribution shifts such that $q_{i}\ll 1$ or $sq_{i}\leq C^{*}$ for some constant $C^{*}>0$, at which point ECR approaches unit success rate while TCR remains bounded by $C/s$, enabling efficient operation with reduced capacity $C=\Theta(1)$ when $C\geq 2C^{*}$.
 Communication Optimization
-The training framework employs the Communication Over Computation (CoC) optimization technique to address performance bottlenecks in LLM training.
-During forward propagation in LLMs, the ColumnParallelLinear and RowParallelLinear components involve sequentially dependent computation (matrix multiplication) and communication (collective operations like AllReduce, AllGather, and ReduceScatter).
-These dependencies lead to inefficient serial execution.
-CoC decomposes these tasks into finer-grained subtasks and merges computation and communication into single kernels, such as MATMUL_ALL_REDUCE and MATMUL_REDUCE_SCATTER, utilizing MTE’s remote memory access capabilities.
-This approach allows for pipeline-style parallel execution and overlapping of computation and communication, significantly enhancing overall efficiency.
+To mitigate the serial execution bottlenecks inherent in LLM training, we implement Communication Over Computation (CoC) optimization, which transforms sequentially dependent matrix multiplication and collective communication operations in parallel linear layers into unified, fine-grained kernels.
+By leveraging MTE’s remote memory access capabilities, CoC fuses computation and communication primitives, enabling pipeline-parallel execution that overlaps previously serialized operations.
 Experiments
 Experimental Setup
-This study employs the Mixtral 8×7B model, incorporating our proposed approach.
-The Mixtral model, comprising 46.7 billion parameters and utilizing Group Query Attention (GQA), features 32 sparse expert blocks with 8 experts in the MoE Feedforward layer, where each token engages the top 2 experts for processing.
-Given the prevalence of long-text corpora in our application scenarios, we extended the sequence length to 32,768 and implemented tailored parallel strategies for cluster scales of 32N, 64N, and 256N, encompassing tensor, pipeline, data, and expert parallelism, with a consistent global batch size of 128.
-For the three cluster scales of 32N, 64N, and 256N, the parallel strategies are set as follows: 32N - tensor parallel (TP=4) / pipeline parallel (PP=4) / data parallel (DP=2) / expert parallel (EP=2), 64N - TP=8 / PP=4 / DP=2 / EP=2, and 256N - TP=8 / PP=8 / DP=4 / EP=2.
-Other details of experimental setup including datasets, environment, and metrics, can be seen in Appendix.
+We implement our approach using Mixtral 8×7B, a 46.7B-parameter model with Group Query Attention (GQA) and 32 sparse MoE blocks.
+Each block contains 8 experts, with tokens routed to their top-2 selections.
+To accommodate long-context applications, we extend the sequence length to 32,768 tokens.
+Our experiments span three cluster scales with tailored parallelization strategies: 32 NPUs (TP=4, PP=4, DP=2, EP=2), 64 NPUs (TP=8, PP=4, DP=2, EP=2), and 256 NPUs (TP=8, PP=8, DP=4, EP=2), maintaining a global batch size of 128 throughout.
+Additional experimental details are provided in the Appendix.
 Efficiency Promotion and Memory Footprint Reduction
-As detailed in Section Method, we consistently use Top-1 routing to ensure the routing implementation aligns with our theoretical framework.
-The Baseline model utilizes a limited expert capacity mode instead of the groupedGEMM scheme, which avoids token dropping, with the capacity factor set to 1.1.
-LocMoE considers data distribution uniformity and estimates expert capacity using a lower bound formula derived from its theoretical conclusions in the first batch, maintaining it as a constant during subsequent training.
-Our approach (abbreviate to ”LocMoE+” in figures) fixes the range of score sums, processes hidden states, and calculates current expert capacity.
-The subsequent analysis addresses the training time, convergence, and memory usage efficiency of these schemes on multiple sizes of Ascend clusters.
+We consistently employ Top-1 routing to align implementation with our theoretical framework.
+The baseline model uses constrained expert capacity rather than groupedGEMM, preventing token dropout with a capacity factor of 1.1.
+LocMoE incorporates distributional uniformity and estimates expert capacity via a theoretically-derived lower bound formula from the initial batch, maintaining this value throughout training.
+Our approach (”LocMoE+” in figures) constrains score sum ranges, processes hidden states, and dynamically calculates expert capacity.
+The subsequent analysis examines training efficiency, convergence, and memory utilization across multiple Ascend cluster configurations.
 Figure 4: The time consumption during training iterations with different schemes and cluster sizes.
-Figure 3a illustrates the time consumption of these methods during the first 1000 iterations of training.
-Due to initialization and some unstable factors, time consumption is recorded starting from the 5th iteration.
-The Baseline model’s time consumption is relatively stable.
-As iterations increase, LocMoE’s time consumption slightly decreases, particularly in 32N and 64N, consistent with the conclusion that locality loss is effective only when the number of experts is greater than or equal to the number of nodes.
-Our approach incurs slightly higher time consumption than LocMoE due to the computational overhead of token rearrangement.
-However, as token features converge, the required tokens gradually decrease and stabilize, leading to a decline in time consumption, which remains stable in subsequent training processes.
-Overall, our approach reduces training time by 2.9% to 13.3% compared to LocMoE, and by 5.4% to 46.6% compared to the Baseline.
-We select 10 iterations at equal intervals from the training iterations to collect data on the time consumption of computation, communication, overlap, and idle periods, as shown in Figure 3(a).
-It is important to note that the data collection operation also introduces some overhead.
-After integrating LocMoE and our approach, the time consumption of each component decreases, with a significantly greater reduction in computation overhead compared to communication overhead.
-Additionally, as the cluster size increases, the proportion of computation/communication overlap decreases, and the magnitude of the reduction in computation overhead diminishes.
-Figure 3(b) illustrates perplexity as a measure of convergence.
-The convergence curves of these approaches indicate normal loss convergence, with our approach not adversely impacting convergence.
+Figure 4 presents the training overhead analysis across the initial 1000 training iterations.
+To ensure measurement stability and exclude initialization artifacts, we commence time profiling from the fifth iteration onward.
+The baseline model exhibits consistent temporal performance throughout the evaluation period.
+In contrast, LocMoE demonstrates a marginal decrease in execution time as training progresses, with this trend being particularly pronounced in the 32N and 64N configurations.
+This observation corroborates our hypothesis that locality-aware optimization achieves optimal efficiency when the number of experts meets or exceeds the number of computational nodes.
+Our proposed method introduces a modest computational overhead relative to LocMoE, attributable to the token rearrangement mechanism.
+However, this overhead diminishes progressively as token representations converge during training.
+Specifically, the convergence of token features leads to a reduction in the number of tokens requiring rearrangement, resulting in stabilized computational costs in later training phases.
+Empirically, our approach achieves a reduction in total training time ranging from 2.9% to 13.3% compared to LocMoE, and from 5.4% to 46.6% relative to the baseline configuration.
+Figure 5 depicts the temporal distribution of computational phases across training.
+We sample performance metrics at ten equidistant intervals throughout training, capturing computation, communication, overlap, and idle time.
+While profiling introduces minor overhead, this methodology provides robust insights into system behavior.
+Both LocMoE and our proposed method demonstrate reduced latency across all components, with computational overhead exhibiting more substantial improvements than communication costs.
+This efficiency gain follows a clear pattern: as cluster size expands, the computation-communication overlap ratio decreases, accompanied by diminishing returns in computational speedup.
+This trend reflects the inherent scalability challenges in distributed MoE architectures.
+Figure 5: The average composition of computation, communication, overlap, and idle with different schemes and cluster sizes.
+Figure 6 validates that these efficiency improvements preserve model quality.
+All methods exhibit comparable convergence trajectories, confirming that our optimization strategy maintains training stability while delivering performance gains.
+The perplexity curves demonstrate that accelerated training does not compromise the fundamental learning dynamics of the model.
+Figure 6: The perplexity during training iterations with different schemes.
+Figure 7 presents the operator-level computational profiling across different hardware components.
+The system leverages AI CORE for matrix multiplication and convolution operations, AI VECTOR CORE for parallelized vector computations, MIX AIC for heterogeneous operator fusion, and AI CPU for specialized AI instruction execution.
+Our token selection strategy yields substantial performance gains: the FFN MatMul operator achieves a 17× speedup compared to the baseline and 2.6× improvement over LocMoE.
+This optimization translates to a 2.8× reduction in cumulative MatMul execution time and a 2.6× decrease in Cube computational load.
+The rearrangement-associated operators (TopK and IndexPutV2) exhibit marginal overhead increases, representing an acceptable trade-off for the significant computational savings achieved through selective token processing.
+Figure 7: The distribution of time consumption for operators.
+Figure 8 analyzes memory consumption patterns during stable training phases, based on 100,000 memory profiling samples per device.
+Our approach demonstrates substantial memory efficiency gains, achieving 4.57-16.27% reduction compared to the baseline and 2.86-10.5% reduction compared to LocMoE.
+The memory optimization exhibits scale-dependent characteristics: larger clusters show reduced computational overhead proportions and correspondingly narrower memory usage differentials.
+Furthermore, our method effectively eliminates transient memory spikes and reduces short-term memory fluctuations, contributing to more predictable and stable resource utilization throughout training.
 Figure 8: print recorded in one acquisition cycle with different schemes and cluster sizes.
-The proportional time consumption at the operator level is depicted in Figure 6.
-Among the components, AI CORE efficiently executes matrix multiplications and convolutions in AI algorithms; AI VECTOR CORE accelerates vector operations through parallel processing; MIX AIC integrates different types of operators and optimizes for multiple tasks; AI CPU is optimized in hardware and instruction sets to better support AI algorithms.
-Our approach selects fewer tokens, resulting in a 17$\times$ performance improvement in the FFN MatMul operator compared to the Baseline and a 2.6$\times$ improvement compared to LocMoE.
-This leads to an overall 2.8$\times$ reduction in the cumulative time consumption of the MatMul operator and a 2.6$\times$ decrease in Cube computing load.
-However, the proportions of TopK and IndexPutV2, related to rearrangement, show a slight increase.
-Figure 6: The distribution of time consumption for operators.
+Expert Homogenization and Load Distribution Analysis
+The Calinski-Harabasz (CH) index (Lima and Cruz 2020) measurements reveal that bidirectional affinity selection significantly enhances token clustering quality in MoE architectures:
+[…]
+While baseline and single-mechanism approaches achieve comparable improvements, the integrated LocMoE+ method demonstrates superior performance, as shown in Figure 9.
+Combining token-to-expert and expert-to-token selection mechanisms creates synergistic effects that accelerate natural clustering tendencies during training.
+Our bidirectional approach establishes a positive feedback loop between token routing and expert affinity, fundamentally enhancing expert utilization efficiency through improved feature organization and specialization.
+Figure 9: The Calinski-Harabasz index across training steps.
+Figure 10 reveals fundamental differences in how various loss functions affect token distribution across experts.
+The baseline approach suffers from severe load imbalance, with certain experts becoming overloaded while others remain idle.
+The auxiliary loss method provides marginal improvements through regularization, yet distribution remains significantly skewed.
+The locality loss demonstrates transformative effects by incorporating architectural topology into the optimization objective, achieving balanced token allocation across all experts through KL divergence constraints that simultaneously minimize inter-node communication and prevent expert collapse.
+Figure 10: The distribution of tokens assigned to experts with different loss function.
+Figure 11 presents the cumulative distribution function (CDF) and empirical cumulative distribution function (ECDF) analysis across these routing methods throughout the training progression.
+The locality loss approach presents distinctly optimal characteristics across both CDF and ECDF measurements, maintaining consistently high performance levels throughout the training process with remarkable stability during later training phases.
+The sustained performance across different probability and load thresholds indicates that incorporating expert-token affinity into the routing objective creates robust optimization dynamics that preserve both routing quality and load distribution efficiency.
+These findings underscore the effectiveness of ETR in addressing the inherent challenges of expert-token assignment optimization, providing a principled foundation for scalable sparse model.
+Figure 11: The CDF and ECDF of different schemes.
+The Performance of Downstream Tasks
+The GDAD benchmark, comprising three distinct evaluation tasks, serves as the primary assessment framework for domain task capabilities, while TeleQnA provides specialized validation for telecommunications knowledge applications.
+Our integrated LocMoE+ approach consistently outperforms both baseline and single-mechanism variants across all evaluation metrics (see Table 1), achieving notable gains through bidirectional selection mechanisms.
+The consistent improvements observed across both general domain tasks and specialized telecommunications applications confirm that the synergistic combination of bidirectional selection mechanisms creates substantial performance advantages while validating the robustness and generalizability of our proposed architecture modifications.
+Table 1: Domain performance promotion obtained by our approach on different datasets.
+GDAD
+GDAD-1 GDAD-2 GDAD-3 Avg TeleQnA
+Baseline 47.8 43.0 65.4 52.8 62.1
+LocMoE (TCR) 55.5 47.6 71.1 59.0 67.6
+LocMoE (ECR) 45.8 45.6 62.8 56.3 61.8
+LocMoE+ 57.4 49.9 74.5 61.5 68.8
+Table 2 presents general performance evaluation results across three widely-recognized benchmarks—MMLU (Hendrycks et al.
+2021) for comprehensive knowledge assessment, GPQA (Rein et al.
+2023) for advanced reasoning capabilities, and HumanEval (Chen et al.
+2021) for code generation proficiency—revealing distinct performance characteristics of our proposed methods.
+The results demonstrate that our bidirectional LocMoE+ approach achieves superior performance in reasoning and coding tasks while maintaining competitive general knowledge capabilities, with individual constraint routing mechanisms exhibiting complementary strengths across different evaluation dimensions.
+While the baseline maintains slight advantage in MMLU, the integrated LocMoE+ approach demonstrates that bidirectional selection mechanisms create meaningful improvements in task-specific capabilities without substantial degradation in general knowledge retention, suggesting that our architectural modifications enhance model specialization for complex reasoning and generation tasks while preserving foundational knowledge capabilities.
+Table 2: General performance comparison of different MoE methods
+Method MMLU GPQA HumanEval
+Baseline 71.8 29.2 40.2
+LocMoE (TCR) 68.4 30.3 52.8
+LocMoE (ECR) 45.8 32.5 57.6
+LocMoE+ 70.4 33.5 67.8
+To enhance conversational capabilities and downstream task adaptability, we conducted supervised fine-tuning on the pre-trained models.
+As shown in Figure 12, our approach demonstrates substantial improvements across multiple evaluation dimensions within the General and Domain-specific Assessment Dataset (GDAD).
+The method achieves an average improvement of approximately 20.1% across 16 sub-capabilities of Domain Task Capability compared to the baseline, with particularly notable gains in rewriting and summary capabilities.
+In the Domain Competency Exam assessments, our approach shows an average improvement of 16% relative to the baseline, with IP Training in digital communications demonstrating the most significant advancement.
+Among the 18 sub-capabilities of General Ability, the method exhibits an improvement of about 13.9% relative to the baseline, with planning capabilities showing the highest enhancement at 26.8%.
 Figure 12: The performance on three categories of GDAD.
-We select a single iteration during the stable training period and describe the per-device memory usage (Allocated) using the first 100,000 samples from its memory monitoring, as shown in Figure 5.
-Overall, our approach achieves memory usage reduction of 4.57% to 16.27% compared to the Baseline and 2.86% to 10.5% compared to LocMoE.
-As cluster size increases, the proportion of computational overhead decreases, and the gap in memory usage narrows.
-Additionally, instantaneous memory peaks gradually disappear, and the fluctuation amplitude of short-term memory also diminishes.
-The Performance of Downstream Tasks
-To enhance the model’s conversational capabilities and adaptability to downstream task, we fine-tuned the pre-trained models.
-As shown in Figure 7, with sufficient supervised fine-tuning (SFT), our approach achieves an average improvement of approximately 20.1% in 16 sub-capabilities of Domain Task Capability, which is a portion of General and Domain-specific Assessment Dataset (GDAD), compared to the Baseline, and an increase of about 3.5% compared to LocMoE.
-The Rewriting and Summary capabilities show the highest improvement, with a 28.2% increase compared to the Baseline and a 6.7% increase compared to LocMoE.
-In the 13 tests of Domain Competency Exam, our approach demonstrates an average improvement of 16% relative to the Baseline and an average increase of approximately 4.8% compared to LocMoE.
-The IP Training in the digital communications domain shows the most significant improvement, with a 27.3% increase compared to the Baseline and a 3.0% increase compared to LocMoE.
-Among the 18 sub-capabilities of General Ability, our approach exhibits an improvement of about 13.9% relative to the Baseline and an average increase of 4.8% compared to LocMoE.
-The capability of Planning demonstrates the highest improvement, with a 26.8% increase compared to the Baseline and a 2.92% increase compared to LocMoE.
-Table 1 presents the holistic evaluation results for multiple datasets, where GDAD-1 represents Domain Task Capability, and the other metrics follow accordingly.
-Notably, due to the 6:4 ratio of Chinese to English data in our incremental pre-training domain data and the 7:3 ratio in the fine-tuning data, our approach achieves an improvement of approximately 13.6% compared to the Baseline and 2.8% compared to LocMoE in the GPQA (Rein et al.
-2023) evaluation, despite the limited data available for training.
-During incremental training and fine-tuning, we incorporated substantial telecommunications domain knowledge, questions, and case studies.
-TeleQnA (Maatouk et al.
-2023), the first benchmark dataset designed to evaluate the knowledge of LLMs in telecommunications, effectively measures the model’s capabilities in this domain.
-Consequently, our approach comprehensively surpasses both the Baseline and LocMoE on this specific dataset.
-Table 1: Performance promotion obtained by our approach on different datasets.
-GDAD
-GDAD-1 GDAD-2 GDAD-3 Avg GPQA TeleQnA
-Baseline 47.8 43.0 65.4 52.8 29.5 62.1
-LocMoE 55.5 47.6 71.1 59.0 32.6 67.6
-LocMoE+ 57.4 49.9 74.5 61.5 33.5 68.8
 Conclusion
-In this paper, we propose a novel expert routing framework that enhances MoE efficiency through three key innovations: an efficient routing mechanism with lightweight computation, a bidirectional expert-token resonance selection mechanism, which combined ECR and TCR, and a dynamic capacity bounds module.
-The framework integrates orthogonal feature extraction and optimized expert localization loss, effectively addressing expert homogeneity while improving routing performance.
-Our local expert strategy demonstrates advantages in both load balancing and communication efficiency.
-Experimental results validate the effectiveness of the proposed framework across multiple benchmarks.
-Our approach achieves performance improvements up to 46.6% (32N) compared to the Baseline and 13.3% (32N) compared to LocMoE, while reducing memory usage by up to 16.27% and 10.5%, respectively.
-To evaluate model performance, all models are evaluated with the open-source datasets GPQA and TeleQnA, and closed domain benchmark GDAD.
-In downstream tasks, our approach outperforms the Baseline by 14.1%, 13.6%, and 9.7% on GDAD, GPQA, and TeleQnA, respectively.
-Future work may explore methods to compress communication data to further reduce communication overhead.
+In this paper, we propose ETR, a fundamentally new approach to MoE routing that solves the longstanding trade-off between computational efficiency and model performance through theoretically-grounded bidirectional selection mechanisms.
+By dynamically coordinating token-choice and expert-choice routing based on training progress, ETR achieves simultaneous improvements in both training efficiency and downstream task quality.
+The substantial performance gains demonstrated across diverse benchmarks, combined with significant reductions in computational overhead, establish ETR as a critical advancement for practical deployment of large-scale sparse models.
+Our theoretical contributions provide new insights into expert-token dynamics that extend beyond incremental optimizations, opening pathways for next-generation MoE architectures.
 Appendix
@@
 Auxiurary Results
-Lemma 8 (Theorem 4 in (Chung and Lu 2006)).
+Lemma 8.
 Let $\bm{X}_{1},\dots,\bm{X}_{n}$ be $n$ independent random variables with
@@
 Then we could derive that
+$\mathcal{P}(\bm{x}\text{ succeed in training})$
 […]
-$=\sum_{i=1}^{n}\mathcal{P}(\bm{o}_{i}\text{ sent to expert }i|\bm{o}_{i}\text{ is in }\bm{x})\cdot\mathcal{P}(\bm{o}_{i}\text{ is in }\bm{x})$
 $=\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$
@@
 Hence, we get the upper bound that
+$\mathcal{P}(\bm{x}\text{ succeed in training})$
 […]
-(Proof.)${{\leq}}\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$ )
 $=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(2nC+\sum_{k=2nC+1}^{s}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\right)$
@@
 Hence, we get the lower bound that
+$\mathcal{P}(\bm{x}\text{ succeed in training})$
 […]
-(Proof.)${{\geq}}\frac{1}{ns}\sum_{i=1}^{n}\sum_{k=1}^{s}p_{i}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})$ )
 $=\frac{1}{ns}\sum_{i=1}^{n}p_{i}\left(\sum_{k=\lceil 1+nC/4\rceil}^{\lfloor 1+nC/2\rfloor}\mathcal{P}(s_{i}<C|\bm{x}_{k}=\bm{o}_{i})\right)$
@@
 By Assumption 4, we derive that $s_{i}\sim\mathcal{B}(s-1,q_{i}),\forall i\in[n]$.
+$\mathcal{P}(\bm{x}\text{ succeed in training})$
 […]
-$=\sum_{i=1}^{n}\mathcal{P}(\text{expert }i\text{ choose }\bm{o}_{i}|\bm{o}_{i}\text{ is in }\bm{x})\mathcal{P}(\bm{o}_{i}\text{ is in }\bm{x})$
 $=\frac{1}{n}\sum_{i=1}^{n}\mathcal{P}(s_{i}\leq C-1,s_{i}\sim\mathcal{B}(s-1,q_{i}))$
@@
 Experimental Setup
-Datasets for Training and Fine-Tuning
+Datasets for Pre-training and Fine-Tuning
 The dataset used in this paper is a self-constructed dataset that integrates knowledge from multiple domains, including wireless, data communication, and cloud-core technologies.
@@
 The domain-specific characteristics include domain knowledge understanding, RAG, FunctionCall, information extraction, multi-turn dialogue, reading comprehension, paraphrasing, and intent recognition.
+The pre-training data comprises 300B tokens in total, with 150B tokens from the ICT domain and 150B tokens sampled from general data.
+The sampling ratios are shown in Table 3.
+For SFT data, we employed a two-stage training: the first stage primarily enhances the model’s logical reasoning capabilities such as multi-task capability, mathematics, puzzle-solving, complex logic, etc.
+The total scale of samples is approximately two million, while the second stage focuses on improving instruction-following abilities, tool function call, sentiment, security, etc.
+The total scale of samples in stage2 is about three million.
+Table 3: Data sources and sampling ratios of general pre-training data.
+Primary Category Secondary Category Tertiary Source Sampling ratio
+General English Webpages Reasoning_steplist 25%
+Model_rewrite 100%
+Books & Papers book3 25%
+bookcorpus 100%
+all_libgen_books 20%
+all_libgen_scihub 10%
+RedPajama_arxiv 25%
+arxiv_latex2Markdown_cleaned 25%
+WebText wiki 100%
+stackexchange_cleaned 20%
+cosmopedia_v2 15%
+General Chinese Webpages aigc_dataset 15%
+Book all_book_deduped 10%
+zh_book_CommonData 10%
+zh_general_STEM 80%
+all_zhiwang 20%
+WebText baike_MBAzhiku_sougou_ye_zhiarge 50%
+baike_sougou_baidu_kuaidong 50%
+wiki 10%
+zhihu_caigou_merged_cleaned 10%
+High-density Knowledge Q&A quiz_data 100%
+Collection density_knowledge 100%
+collection_updated 100%
+english_question_and_answer 100%
+annealing 100%
+Code Code code_python_edu_high_quality 30%
+Forum CSDN 20%
+Ultra_textbooks 100%
 Experimental Environment
@@
 Every 8 NPUs are mounted on the same Atlas 800T A2 server, which internally adopts a fullmesh networking scheme, meaning that any two NPUs are interconnected.
-The version of the Ascend Hardware Development Kit (HDK) is 23.0.2.1, and the version of the Compute Architecture for Neural Networks (CANN) suite is 7.0.0, which is the commercial release version for Q4 2023.
-The models in this paper use ModelLink, an LLM training framework based on the Ascend architecture, and run in the torch_npu 5.0.0 environment.
-Model Configuration
-This paper adopts the Mixtral 8$\times$7B model with an MoE structure, released in December 2023, as the backbone and embeds LocMoE and our approach.
-Mixtral has 46.7B parameters and uses the Group Query Attention (GQA) mechanism to compute attention.
-It contains 32 sparse expert blocks, with the MoE Feedforward layer comprising 8 experts, and each token selects the top 2 experts for processing.
-In the application scenarios of this paper, the corpora are generally long texts; thus, the sequence length is adjusted to 32768.
-For the three cluster scales of 32N, 64N, and 256N, the parallel strategies are set as follows: 32N - tensor parallel (TP=4) / pipeline parallel (PP=4) / data parallel (DP=2) / expert parallel (EP=2), 64N - TP=8 / PP=4 / DP=2 / EP=2, and 256N - TP=8 / PP=8 / DP=4 / EP=2.
-The batch size is set to 128.
 Evaluation Metrics and Datasets
@@
 For subjective questions, the cosine similarity between the model output and the standard answer is used as the score.
-In addition, this paper also employs C-Eval (Huang et al.
-2024b) and TeleQnA (Maatouk et al.
+In addition, this paper also employs GPQA (Rein et al.
+2023) and TeleQnA (Maatouk et al.
 2023) to evaluate the model’s Chinese language capability.